### v1

# MTVQA: Benchmarking Multilingual Text-Centric Visual Question Answering

Abstract
Text-Centric Visual Question Answering (TEC-VQA) in its proper format not only facilitates human-machine interaction in text-centric visual environments but also serves as a de facto gold proxy to evaluate AI models in the domain of text-centric scene understanding. However, most TEC-VQA benchmarks have focused on high-resource languages like English and Chinese. Despite pioneering works to expand multilingual QA pairs in non-text-centric VQA datasets using translation engines, the translation-based protocol encounters a substantial “Visual-textual misalignment” problem when applied to TEC-VQA. Specifically, it prioritizes the text in question-answer pairs while disregarding the visual text present in images. Furthermore, it does not adequately tackle challenges related to nuanced meaning, contextual distortion, language bias, and question-type diversity. In this work, we address the task of multilingual TEC-VQA and provide a benchmark with high-quality human expert annotations in 9 diverse languages, called MTVQA. To our knowledge, MTVQA is the first multilingual TEC-VQA benchmark to provide human expert annotations for text-centric scenarios. Further, by evaluating several state-of-the-art Multimodal Large Language Models (MLLMs), including GPT-4V, on our MTVQA dataset, it is evident that there is still room for performance improvement, underscoring the value of our dataset. We hope this dataset will provide researchers with fresh perspectives and inspiration within the community.
footnotetext: Equal contribution. 🖂Corresponding author.
1 Introduction
In the era of burgeoning AI, especially in LLMs/MLLMs [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17], Text-Centric Visual Question Answering (TEC-VQA[18, 19, 10, 9, 13, 12, 11] has served as a de facto gold proxy to evaluate AI models in the domain of text-centric scene understanding. Compared with general VQA [18, 20, 21, 19, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33], TEC-VQA places greater emphasis on answering questions that require understanding textual information within images. It provides a streamlined avenue for individuals without specialized expertise to articulate their requirements and access applications in text-centric visual environments. However, the majority of advancements in TEC-VQA have predominantly concentrated on high-resource languages, e.g., English [18, 19, 20, 23], Chinese [34, 31], Japanese [35, 36] and etc., thus restricting the applicability of AI models to the global community, particularly populations speaking low-resource languages.
To tackle the problem of language diversity, several seminal studies [37, 38, 39] in the general VQA field, leverage off-the-shelf translation engines to expand existing question-answer pairs from high-resource languages to their multilingual counterparts including low-resource ones. However, when applied to TEC-VQA, this translation-based approach may fall prey to the “Visual-textual misalignment” problem as only the text in question-answer pairs can be processed, while the visual text present in the images is overlooked. Not to mention issues such as nuanced meaning, contextual distortion, language bias, and question type diversity further render the transferability of the translation protocol infeasible for TEC-VQA. The status quo begs for a question: “How can we address the visual-textual misalignment problem for multilingual TEC-VQA and what we stand in the MLLM era?
Refer to caption Refer to caption
Figure 1: Left: overview of various categories of text-centric images. Right: distribution of image and QA pairs over the 9 languages in MTVQA benchmark.
Figure 2: Multilingual VQA examples selected from four languages. From left to right: Arabic (AR), Russian (RU), Thai (TH), Vietnamese (VI). The corresponding translations in English are in brackets.
In this work, to answer the question above, we establish MTVQA, a novel and high-quality multilingual TEC-VQA benchmark, where all images are collected from real-world and meticulously annotated by human experts in nine languages: Arabic (AR), Korean (KO), Japanese (JA), Thai (TH), Vietnamese (VI), Russian (RU), French (FR), German (DE), and Italian (IT). More concretely, to ensure the visual-textual alignment at most, the annotation process follows the raise-then-correct paradigm, where a group of human annotators raises several distinct questions, ranging from simple content extraction to text-related reasoning, and subsequently provides answers. These QA pairs are then double-checked by another group to ensure accuracy and consistency. Consequently, as illustrated in Fig. 5, 6,678 training images and 21,829 question-answer pairs, as well as 2,116 test images and 6,778 question-answer pairs are obtained, covering several fine-grained scenarios, such as menus, logos, maps, bills, PPTs, research papers, and etc. To our best knowledge, MTVQA is the first TEC-VQA dataset to provide native human annotations for multilingual text-rich scenarios, especially for low-source languages. Furthermore, we investigate recent representative MLLMs, including GPT-4V, Gemini, QwenVL etc., by juxtaposing experimental results regarding their performance on our newly proposed MTVQA. Both for general MLLMs and document-focused ones, the results unequivocally demonstrate that opportunities for improvement persist within these MLLMs when applied in multilingual text-rich scenarios.
In summary, the main contributions of this paper can be categorized into three points:
  • We introduce the MTVQA dataset, to the best of our knowledge, which is the first multilingual TEC-VQA benchmark to provide human expert annotations for text-centric scenarios.
  • We benchmark the state-of-the-art MLLMs on our new dataset and show there is still room for performance improvement for these models under multilingual text-rich scenarios.
  • We propose a set of baselines for multilingual TEC-VQA tasks.
2 Related Work
2.1 LLMs/MLLMs for text-centric VQA
Recent advancements in LLMs/MLLMs [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17] have revolutionized VQA tasks, as demonstrated by the remarkable zero-shot performance of these models. Notably, the high generalizability of LLMs/MLLMs, when explicitly trained on visual text understanding datasets and fine-tuned with instructions, has significantly enhanced their application in text-centric VQA scenarios [10, 9, 13, 12, 11]. For example, LLaVAR [40], UniDoc [10], which extend LLaVA [30] into the realm of document understanding, pioneering the text-centric VQA of MLLMs by training them to predict texts and coordinates from document images. Furthermore, DocPedia [9] operates visual input in the frequency domain rather than in space, which enables higher input resolution without increasing the input sequence. Lately, mPLUG-DocOwl [41], Qwen-VL [6], and TextMonkey [12] leverage publicly available document-related VQA datasets to further enhance the text-centric VQA capabilities. Despite the promising results achieved by existing LLMs/MLLMs in text-centric VQA tasks, their focus on high-resource languages such as English or Chinese has posed challenges in achieving reasonable performance for low-resource languages. This is primarily due to the lack of data or benchmarks for these low-resource languages.
2.2 Multilingual text-centric VQA Benchmarks
VQA has garnered significant attention in recent years, with numerous studies, datasets, and benchmarks being proposed to advance the field [18, 20, 21, 19, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33]. Many datasets have been created that encompass scene text of various domains, including natural images [18, 19], scanned documents [20, 23], book and movie covers [22]. One notable limitation of these datasets is their predominant focus on English [18, 19, 20, 23] or other high-resource languages such as Chinese [34, 31] and Japanese [35, 36], which restricts the applicability of VQA systems for low-resource languages such as Thai and Vietnamese.
There is a recent effort toward extending VQA tasks to a wider range of languages [42, 38, 43, 39, 44, 37] by providing a multilingual VQA datasets. For example,  Gao et al. [31] created a free-form bilingual VQA dataset (FM-IQA) contains over 150,000 images and 310,000 freestyle Chinese question-answer pairs and their English translations. Raj Khan et al. [37] developed a large-scale multilingual and code-mixed VQA dataset (MuCo-VQA) supporting five languages. Of more relevance are the works xGQA (8 languages) [38] and MaXM (7 languages) [39], which apply translation-based protocols to expand VQA data beyond English. However, the translation-based multilingual VQA datasets inherently face issues, such as the “Visual-textual misalignment” problem, where only the text in question-answer pairs is processed, while the visual text in images is overlooked. Additionally, the nuanced meaning and context are often distorted; language bias introduced by machine translation models, and the coverage of certain question types is limited, as highlighted by Changpinyo et al. [39]. Moreover, none of the previous multilingual datasets focus on text-centric scenarios where multilingual text frequently occurs.
Our benchmark distinguishes itself by focusing on multilingual text-centric VQA scenarios using human expert annotations. To the best of our knowledge, the MTVQA benchmark is the first dataset to provide native human annotations for such scenarios. It covers 9 languages, thereby facilitating the training and evaluation of multilingual models in diverse linguistic contexts. Additionally, our dataset can gauge the VQA system’s ability for not only high-resource languages but also those that are typically underrepresented in current datasets [18, 19, 20, 23, 31].
The MTVQA benchmark addresses a significant gap in existing datasets by catering to the crucial needs of low-resource languages through annotations from native speakers across multiple languages. Our pioneering efforts distinctly position the MTVQA benchmark as a unique multilingual VQA resource, advancing the frontier of machine learning research.
3 MTVQA Benchmark
The MTVQA Benchmark covers 9 languages: Arabic (AR), Korean (KO), Japanese (JA), Thai (TH), Vietnamese (VI), Russian (RU), French (FR), German (DE), and Italian (IT). In this section, we describe in detail how we establish the MTVQA benchmark, including the collection of raw image data and two-round human expert annotations, which are independent of each other.
3.1 Data Collection
Our purpose is to develop a multilingual VQA benchmark capable of evaluating the QA performance of MLLMs in multilingual text-centric scenarios, thus the raw data collection process is mainly oriented towards text-centric images from natural scenarios and document scenarios. To ensure the diversity and quality of data, we collect not only the raw image data from publicly available datasets, including the multilingual scene text recognition images from MLT2019 [45] and PowerPoint slides (PPTs) sourced from the internet, but also the data from countries of each language. Furthermore, the collected data includes multiple fine-grained scenarios (Fig. 1), such as menus, logos, maps, bills, PPTs, research papers, and etc. As a result, we gather a total of 1,220 images from document scenarios and 876 images from natural scenarios in the test set of the MTVQA benchmark. To ensure the visual-textual alignment, for text-rich images lacking text and language annotations, we subject them to a standardized data cleaning process, which includes text recognition and language classification. Afterward, we organize all the text-rich images we have obtained into language-specific groups, preparing them for the subsequent stage of data annotation.
3.2 Human Expert Annotation
In order to obtain informative and accurate text-related QA pairs on the language-specific grouped images, we recruit a group of annotators with expertise from local regions of each language. It is worth noting that all these annotators are native speakers of their respective languages, ensuring their deep understanding and proficiency in the linguistic nuances and cultural context necessary for precise annotations. Considering the subjective nature of the text-image understanding task, we have implemented a further division within the annotation team. This division involves separating the team into two independent groups, with one group dedicated to generating and responding to questions based on the provided images, while the other group focuses on evaluating and correcting the QA pair results. This raise-then-correct paradigm ensures a comprehensive and reliable assessment of the text-image understanding process. Additionally, each language’s annotation results undergo a 10% sampling inspection by a quality inspector. If the QA pairs fail to meet the criteria, they are sent back for re-annotation. Prior to commencing the formal human expert annotation task, all annotators undergo unified training and receive annotation examples. The brief diagram of the two-round annotation process is shown in Figure 3 and we elaborate on it in the following subsections.
Table 1: Mean lengths of question-answer pairs in different languages of training set and test set, using GPT-4o tokenizer.
AR DE FR IT JA KO RU TH VI
Training Set
Question 8.29 8.72 9.73 12.05 12.43 11.74 11.56 11.35 11.21
Answer 9.66 6.96 7.34 11.24 12.70 13.56 12.00 11.26 13.31
Test Set
Question 8.08 8.29 9.76 11.93 12.48 12.2 11.65 10.98 10.99
Answer 7.95 6.67 6.61 11.04 12.55 13.61 14.42 12.08 13.00
First Round Questioning and Answering. For the first round of annotation tasks, we assigned 3 annotators for each language to manually generate original QA results. Given a text-centric image from our collection, annotators are first required to read the texts in the image and analyze other contents in the image in a comprehensive and detailed manner. They must then raise 4 meaningful and distinct questions based on the content in the image and give the answers. All annotators adhere to the following criteria: (1) the first three questions should satisfy that answering these questions requires direct reading of the textual information in the image, (2) the fourth question requires reasoning about the text in the image to answer (3) the questions and answers must be reasonably correct and consistent with the content of the image, and (4) the answer should be as concise as possible and free of nonsense (e.g., when the question is “When is the volunteer recruitment period”, the answer should be “9:00-16:00” rather than “The volunteer recruitment period is 9:00-16:00”). It’s worth mentioning that our requirement for concise answers is to make the evaluation process more friendly and more reliable, cause we try to keep the evaluation metrics unaffected by extraneous content in the answer sentence.
Second round Evaluation and Correction. To reduce the effect of human subjective cognitive bias on our MTVQA benchmark and get high-quality question-answer pairs, we assigned 2 annotators for each language for the annotation evaluation and correction process. Based on the provided images and the first-round annotation results, the annotators must follow these rules of judgment and steps for the annotation: (1) Whether the question is related to the text in the image. If not, discard the current question-answer pair, (2) Whether the answer is correct. If not, modify the answer, and (3) Whether the answer repeats the content from the question. If so, remove the repeated content to ensure a concise answer.
Figure 3: A brief diagram of the annotation process.
3.3 Data Statistics
We instruct the annotators to complete the above human expert annotation work towards the text-centric VQA tasks and construct the MTVQA benchmark consisting of 8,794 images and 28,607 question-answer pairs that cover the 9 languages. The MTVQA benchmark is divided into a training set containing 6,678 images and 21,829 question-answer pairs, and a test set containing 2,116 images and 6,778 question-answer pairs. The detailed data distribution can be seen in Figure 1. To visualize the vocabulary richness of our benchmark, we calculate the word frequencies for each language and present them in the form of word clouds as shown in Figure 4. In Figure 5 we demonstrate the statistics of the question and answer lengths using GPT-4o tokenizer.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 4: Top answer word clouds of different languages.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 5: Statistics of question and answer lengths of different languages aggregating training and test set, using GPT-4o tokenizer.
4 Experiments
4.1 Baseline Models
For the MTVQA benchmark, we evaluate the following instruction-tuned general MLLMs, (1) Open-source MLLMs: InternVL-V1.5 [46], InternLM-Xcomposer2-4KHD [15], Mini-Gemini-HD-34B [16], Llava-Next-34B [17], DeepSeek-VL [7], YI-VL-34B [8], TextSquare [13], TextMonkey [12] and mPLUG-DocOwl 1.5 [11]; (2) Closed-source MLLMs: GPT-4V, Gemini Ultra, QwenVL Max, QwenVL Plus, Claude3 Opus, Claude3 Sonnet and GLM4V. For the closed-source MLLMs, we use the chat version through the official APIs, while for the open-source MLLMs, we utilize the instruct versions. It is noted that all the model weights of the open-source MLLMs evaluated in our experiments could be downloaded from the HuggingFace Model Hub. For the open-source MLLMs, the model size varies from 7b to 34b.
4.2 Implementation Details
We conduct the evaluation experiments over the baseline MLLMs with their default settings, ignoring the effect of generation configuration on the results. To make the output of MLLMs more evaluation-friendly, we design the following prompt format to limit the output length: “Answer the question using a word or phrase in the language of the question. + <Question>”, where <Question> represents the corresponding question of the input image. The extra prefixes added to the raw question could limit the answer to be as concise as possible. Besides, we utilize the InternLM-Xcomposer2-4KHD [15] as the backbone for the instructional fine-tuning experiment on the MTVQA training set. In the instructional fine-tuning process, we follow the default training settings [15] with “HD-16” and train on MTVQA training set for 1 epoch.
4.3 Evaluation Results
[…]
Table 2: Performance of the leading closed-source and open-source MLLMs on the MTVQA benchmark.
AR DE FR IT JA KO RU TH VI Avg.
Closed-Source MLLMs
GPT-4V 11.5 31.5 40.4 32.3 11.5 16.7 10.3 15.0 28.9 22.0
Gemini Ultra 14.7 32.3 40.0 31.8 12.3 17.2 11.8 20.3 28.6 23.2
QwenVL Max 7.7 31.4 37.6 30.2 18.6 25.4 10.4 4.8 23.5 21.1
QwenVL Plus 4.8 28.8 33.7 27.1 12.8 19.9 9.4 5.6 18.1 17.8
Claude3 Opus 15.1 33.4 40.6 34.4 19.4 27.2 13.0 19.5 29.1 25.7
Claude3 Sonnet 10.5 28.9 35.6 31.8 13.9 22.2 11.0 15.2 20.8 21.1
GLM4V 0.3 30.0 34.1 30.1 3.4 5.7 3.0 3.5 12.3 13.6
Open-Source MLLMs
InternVL-V1.5 [46] 3.4 27.1 31.4 27.1 9.9 9.0 4.9 8.7 12.4 14.9
Mini-Gemini-HD-34B [16] 2.2 25.0 29.2 25.5 6.1 8.6 4.1 4.3 11.8 13.0
Llava-Next-34B [17] 3.3 24.0 28.0 22.3 3.6 6.1 2.6 0.4 9.8 11.1
DeepSeek-VL [7] 0.6 14.2 15.3 15.2 2.9 3.8 1.6 0.9 5.2 6.6
YI-VL-34B [8] 1.7 13.5 15.7 12.1 4.8 5.2 0.8 3.5 4.1 6.8
TextSquare [13] 3.7 27.0 30.8 26.7 3.2 7.2 6.7 5.2 12.4 13.6
TextMonkey [12] 2.0 18.1 19.9 22.1 4.6 7.2 3.2 0.9 11.1 9.9
mPLUG-DocOwl 1.5 [11] 1.0 13.9 14.9 18.2 2.9 5.0 2.0 0.9 6.4 7.2
Xcomposer2-4KHD [15] 2.0 20.6 23.2 21.6 5.6 7.7 4.1 6.1 10.1 11.2
Xcomposer-SFT 11.8 31.7 37.4 29.3 14.5 12.9 5.8 13.9 20.2 19.7
5 Limitation
The current iteration of MTVQA exhibits certain constraints that warrant attention. Primarily, the linguistic diversity incorporated is not exhaustive; several lesser-spoken languages remain unrepresented. Future enhancements will aim to broaden the multilingual scope of the dataset. Additionally, the dataset currently offers a singular canonical response for each question. Recognizing the multifaceted nature of the inquiry, subsequent versions will endeavor to include a spectrum of plausible answers to reflect the varied perspectives inherent to each question.
6 Conclusion
In this paper, we introduce MTVQA, a multilingual TEC-VQA benchmark featuring high-quality human expert annotations in 9 diverse languages. We believe that MTVQA is the first benchmark of its kind to provide fully manual annotations specifically tailored to text-centric scenarios. The results obtained from both closed- and open-source MLLMs on our MTVQA dataset indicate that there is still room for improving their performance in multilingual text-centric scenarios. Although the current version of MTVQA has constraints regarding linguistic diversity and singular responses per question, we are confident that this dataset can still inspire researchers within the TEC-VQA community with new perspectives and ideas.

### v2

@@
 Text-Centric Visual Question Answering (TEC-VQA) in its proper format not only facilitates human-machine interaction in text-centric visual environments but also serves as a de facto gold proxy to evaluate AI models in the domain of text-centric scene understanding.
-However, most TEC-VQA benchmarks have focused on high-resource languages like English and Chinese.
-Despite pioneering works to expand multilingual QA pairs in non-text-centric VQA datasets using translation engines, the translation-based protocol encounters a substantial “Visual-textual misalignment” problem when applied to TEC-VQA.
+Nonetheless, most existing TEC-VQA benchmarks have focused on high-resource languages like English and Chinese.
+Despite pioneering works to expand multilingual QA pairs in non-text-centric VQA datasets through translation engines, the translation-based protocol encounters a substantial “visual-textual misalignment” problem when applied to TEC-VQA.
 Specifically, it prioritizes the text in question-answer pairs while disregarding the visual text present in images.
-Furthermore, it does not adequately tackle challenges related to nuanced meaning, contextual distortion, language bias, and question-type diversity.
-In this work, we address the task of multilingual TEC-VQA and provide a benchmark with high-quality human expert annotations in 9 diverse languages, called MTVQA.
-To our knowledge, MTVQA is the first multilingual TEC-VQA benchmark to provide human expert annotations for text-centric scenarios.
-Further, by evaluating several state-of-the-art Multimodal Large Language Models (MLLMs), including GPT-4V, on our MTVQA dataset, it is evident that there is still room for performance improvement, underscoring the value of our dataset.
-We hope this dataset will provide researchers with fresh perspectives and inspiration within the community.
-footnotetext: Equal contribution. 🖂Corresponding author.
+Moreover, it fails to address complexities related to nuanced meaning, contextual distortion, language bias, and question-type diversity.
+In this work, we tackle multilingual TEC-VQA by introducing MTVQA, the first benchmark featuring high-quality human expert annotations across 9 diverse languages, consisting of 6,778 question-answer pairs across 2,116 images.
+Further, by comprehensively evaluating numerous state-of-the-art Multimodal Large Language Models (MLLMs), including GPT-4o, GPT-4V, Claude3, and Gemini, on the MTVQA dataset, it is evident that there is still a large room for performance improvement, underscoring the value of MTVQA.
+Additionally, we supply multilingual training data within the MTVQA dataset, demonstrating that straightforward fine-tuning with this data can substantially enhance multilingual TEC-VQA performance.
+We aspire that MTVQA will offer the research community fresh insights and stimulate further exploration in multilingual visual text comprehension.
+The project homepage is at https://bytedance.github.io/MTVQA/.
+footnotetext: Equal contribution. 🖂 Corresponding authors.
 1 Introduction
-In the era of burgeoning AI, especially in LLMs/MLLMs [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17], Text-Centric Visual Question Answering (TEC-VQA[18, 19, 10, 9, 13, 12, 11] has served as a de facto gold proxy to evaluate AI models in the domain of text-centric scene understanding.
+In the era of burgeoning AI, especially in MLLMs [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18], Text-Centric Visual Question Answering (TEC-VQA[19, 20, 11, 10, 14, 13, 12] has served as a de facto gold proxy to evaluate AI models in the domain of text-centric scene understanding.
 Compared with general VQA [19, 21, 22, 20, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34], TEC-VQA places greater emphasis on answering questions that require understanding textual information within images.
@@
 However, when applied to TEC-VQA, this translation-based approach may fall prey to the “visual-textual misalignment” problem as only the text in question-answer pairs can be processed, while the visual text present in the images is overlooked.
-Not to mention issues such as nuanced meaning, contextual distortion, language bias, and question type diversity further render the transferability of the translation protocol infeasible for TEC-VQA.
-The status quo begs for a question: “How can we address the visual-textual misalignment problem for multilingual TEC-VQA and what we stand in the MLLM era?
+Although the visual text can be obtained by an OCR (Optical Character Recognition) engine and then translated into English, this indirect process could introduce issues such as nuanced meaning, contextual distortion, language bias, and question type diversity.
+These factors further render the transferability of the translation protocol infeasible for TEC-VQA.
+Taking the second case in Fig.
+1 as an example, if either the visual text or the question in Russian is recognized or translated problematically, then the question would never be answered.
+The status quo begs for a question: “Can we address the visual-textual misalignment problem directly from the source language per se for multilingual TEC-VQA and what we stand in the MLLM era?
+Figure 1: Multilingual text-centric VQA visualization selected from four languages.
+From left to right: Arabic (AR), Russian (RU), Thai (TH), Vietnamese (VI).
+The corresponding translations in English are in brackets.
+More examples see Figure 7.
 Refer to caption Refer to caption
-Figure 1: Left: overview of various categories of text-centric images. Right: distribution of image and QA pairs over the 9 languages in MTVQA benchmark.
-Figure 2: Multilingual VQA examples selected from four languages. From left to right: Arabic (AR), Russian (RU), Thai (TH), Vietnamese (VI). The corresponding translations in English are in brackets.
+Figure 2: Left: overview of various categories of text-rich images. Right: image and QA pairs distribution over the 9 languages in MTVQA benchmark.
 In this work, to answer the question above, we establish MTVQA, a novel and high-quality multilingual TEC-VQA benchmark, where all images are collected from real-world and meticulously annotated by human experts in nine languages: Arabic (AR), Korean (KO), Japanese (JA), Thai (TH), Vietnamese (VI), Russian (RU), French (FR), German (DE), and Italian (IT).
 More concretely, to ensure the visual-textual alignment at most, the annotation process follows the raise-then-correct paradigm, where a group of human annotators raises several distinct questions, ranging from simple content extraction to text-related reasoning, and subsequently provides answers.
-These QA pairs are then double-checked by another group to ensure accuracy and consistency.
+Another group then double-checks these QA pairs to ensure accuracy and consistency.
 Consequently, as illustrated in Fig.
-5, 6,678 training images and 21,829 question-answer pairs, as well as 2,116 test images and 6,778 question-answer pairs are obtained, covering several fine-grained scenarios, such as menus, logos, maps, bills, PPTs, research papers, and etc.
-To our best knowledge, MTVQA is the first TEC-VQA dataset to provide native human annotations for multilingual text-rich scenarios, especially for low-source languages.
-Furthermore, we investigate recent representative MLLMs, including GPT-4V, Gemini, QwenVL etc., by juxtaposing experimental results regarding their performance on our newly proposed MTVQA.
-Both for general MLLMs and document-focused ones, the results unequivocally demonstrate that opportunities for improvement persist within these MLLMs when applied in multilingual text-rich scenarios.
+2, 6,678 training images and 21,829 question-answer pairs, as well as 2,116 test images and 6,778 question-answer pairs are obtained, covering more than 20 fine-grained scenarios from both documents and natural scenes, such as menus, logos, maps, bills, PPTs, research papers, and etc.
+To our knowledge, MTVQA is the first TEC-VQA dataset to provide native human annotations for multilingual text-rich scenarios, especially for low-source languages.
+Figure 3: Left: comparison of the overall performance of various MLLMs in the MTVQA benchmark. Right: comparison of the performance exhibited by MLLMs in the 9 languages of the MTVQA.
+We further investigate recent representative MLLMs in Fig.
+3, including GPT-4o, GPT-4V, Gemini, QwenVL, etc., and their performance on our proposed MTVQA.
+As vividly revealed in Fig.
+3 (a), GPT-4o is the top performer, followed by Claude3 Opus and Gemini Ultra, with closed-source MLLMs generally leading.
+General MLLMs like QwenVL Max and QwenVL Plus show mid-tier performance, while text-centric MLLMs lag behind.
+Fig.
+3(b) shows closed-source MLLMs consistently outperforming others, especially in French (FR) and German (DE), while languages like Arabic (AR) and Thai (TH) witness lower scores across all categories.
+The results unequivocally demonstrate that opportunities for improvement persist within existing MLLMs when applied in multilingual text-rich scenarios.
 In summary, the main contributions of this paper can be categorized into three points:
-  • We introduce the MTVQA dataset, to the best of our knowledge, which is the first multilingual TEC-VQA benchmark to provide human expert annotations for text-centric scenarios.
-  • We benchmark the state-of-the-art MLLMs on our new dataset and show there is still room for performance improvement for these models under multilingual text-rich scenarios.
-  • We propose a set of baselines for multilingual TEC-VQA tasks.
+  • We coin the MTVQA dataset, which to the best of our knowledge is the first multilingual TEC-VQA benchmark providing human expert annotations to solve “visual-textual misalignment” problem in multilingual text-centric scenarios.
+  • We benchmark the state-of-the-art MLLMs on our new dataset and show that there are still opportunities for improvement on even the most advanced MLLMs when it comes to multilingual text-rich scenarios.
+  • We establish a set of new multilingual TEC-VQA baselines for both closed-source and general-purpose, text-centric MLLMs.
 2 Related Work
-2.1 LLMs/MLLMs for text-centric VQA
-Recent advancements in LLMs/MLLMs [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17] have revolutionized VQA tasks, as demonstrated by the remarkable zero-shot performance of these models.
-Notably, the high generalizability of LLMs/MLLMs, when explicitly trained on visual text understanding datasets and fine-tuned with instructions, has significantly enhanced their application in text-centric VQA scenarios [10, 9, 13, 12, 11].
+2.1 MLLMs for Text-centric VQA
+Recent advancements in MLLMs [2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18] have revolutionized VQA tasks, as demonstrated by the remarkable zero-shot performance of these models.
+Notably, the high generalizability of MLLMs, when explicitly trained on visual text understanding datasets and fine-tuned with instructions, has significantly enhanced their application in text-centric VQA scenarios [11, 10, 14, 13, 12].
 For example, LLaVAR [41], UniDoc [11], which extend LLaVA [31] into the realm of document understanding, pioneering the text-centric VQA of MLLMs by training them to predict texts and coordinates from document images.
@@
 Lately, mPLUG-DocOwl [42], Qwen-VL [7], and TextMonkey [13] leverage publicly available document-related VQA datasets to further enhance the text-centric VQA capabilities.
-Despite the promising results achieved by existing LLMs/MLLMs in text-centric VQA tasks, their focus on high-resource languages such as English or Chinese has posed challenges in achieving reasonable performance for low-resource languages.
+Despite the promising results achieved by existing MLLMs in text-centric VQA tasks, their focus on high-resource languages such as English and Chinese has posed challenges in achieving reasonable performance for low-resource languages.
 This is primarily due to the lack of data or benchmarks for these low-resource languages.
@@
 One notable limitation of these datasets is their predominant focus on English [19, 20, 21, 24] or other high-resource languages such as Chinese [35, 32] and Japanese [36, 37], which restricts the applicability of VQA systems for low-resource languages such as Thai and Vietnamese.
-There is a recent effort toward extending VQA tasks to a wider range of languages [42, 38, 43, 39, 44, 37] by providing a multilingual VQA datasets.
+There are some recent efforts toward extending VQA tasks to a wider range of languages [43, 39, 44, 40, 45, 38] by providing a multilingual VQA datasets.
 For example,  Gao et al.
-[31] created a free-form bilingual VQA dataset (FM-IQA) contains over 150,000 images and 310,000 freestyle Chinese question-answer pairs and their English translations.
+[32] created a free-form bilingual VQA dataset (FM-IQA) containing over 150,000 images and 310,000 freestyle Chinese question-answer pairs and English translations.
 Raj Khan et al.
@@
 Of more relevance are the works xGQA (8 languages) [39] and MaXM (7 languages) [40], which apply translation-based protocols to expand VQA data beyond English.
-However, the translation-based multilingual VQA datasets inherently face issues, such as the “Visual-textual misalignment” problem, where only the text in question-answer pairs is processed, while the visual text in images is overlooked.
-Additionally, the nuanced meaning and context are often distorted; language bias introduced by machine translation models, and the coverage of certain question types is limited, as highlighted by Changpinyo et al.
-[39].
+However, the translation-based multilingual VQA datasets inherently face issues, such as the “visual-textual misalignment” problem, where only the textual information in question-answer pairs is considered, while the visual text in images is overlooked.
+Additionally, the nuanced meaning and context are often distorted; language bias is introduced by machine translation models, and the coverage of certain question types is limited, as highlighted by [40].
 Moreover, none of the previous multilingual datasets focus on text-centric scenarios where multilingual text frequently occurs.
-Our benchmark distinguishes itself by focusing on multilingual text-centric VQA scenarios using human expert annotations.
-To the best of our knowledge, the MTVQA benchmark is the first dataset to provide native human annotations for such scenarios.
+Our benchmark, MTVQA, distinguishes itself by focusing on multilingual text-centric VQA scenarios using human expert annotations.
 It covers 9 languages, thereby facilitating the training and evaluation of multilingual models in diverse linguistic contexts.
 Additionally, our dataset can gauge the VQA system’s ability for not only high-resource languages but also those that are typically underrepresented in current datasets [19, 20, 21, 24, 32].
-The MTVQA benchmark addresses a significant gap in existing datasets by catering to the crucial needs of low-resource languages through annotations from native speakers across multiple languages.
-Our pioneering efforts distinctly position the MTVQA benchmark as a unique multilingual VQA resource, advancing the frontier of machine learning research.
 3 MTVQA Benchmark
-The MTVQA Benchmark covers 9 languages: Arabic (AR), Korean (KO), Japanese (JA), Thai (TH), Vietnamese (VI), Russian (RU), French (FR), German (DE), and Italian (IT).
-In this section, we describe in detail how we establish the MTVQA benchmark, including the collection of raw image data and two-round human expert annotations, which are independent of each other.
+The MTVQA benchmark is meticulously established to evaluate the multilingual text comprehension performance of Multimodal Large Language Models (MLLMs).
+MTVQA covers nine languages: Arabic (AR), Korean (KO), Japanese (JA), Thai (TH), Vietnamese (VI), Russian (RU), French (FR), German (DE), and Italian (IT).
+The data construction process involves a comprehensive collection of text-rich images and a two-round human expert annotation process.
+Figure 4: A brief diagram of the annotation process.
 3.1 Data Collection
-Our purpose is to develop a multilingual VQA benchmark capable of evaluating the QA performance of MLLMs in multilingual text-centric scenarios, thus the raw data collection process is mainly oriented towards text-centric images from natural scenarios and document scenarios.
-To ensure the diversity and quality of data, we collect not only the raw image data from publicly available datasets, including the multilingual scene text recognition images from MLT2019 [45] and PowerPoint slides (PPTs) sourced from the internet, but also the data from countries of each language.
-Furthermore, the collected data includes multiple fine-grained scenarios (Fig.
-1), such as menus, logos, maps, bills, PPTs, research papers, and etc.
-As a result, we gather a total of 1,220 images from document scenarios and 876 images from natural scenarios in the test set of the MTVQA benchmark.
-To ensure the visual-textual alignment, for text-rich images lacking text and language annotations, we subject them to a standardized data cleaning process, which includes text recognition and language classification.
-Afterward, we organize all the text-rich images we have obtained into language-specific groups, preparing them for the subsequent stage of data annotation.
+The raw data collection is oriented towards gathering text-rich images from various natural and document scenarios, ensuring diversity and quality.
+This includes images from publicly available datasets(e.g., ICDAR MLT19 [46]) and those sourced from the internet, such as menus, logos, maps, bills, PowerPoint slides (PPTs), research papers, etc, as shown in Fig.
+2.
+A total of 1,220 images from document scenarios and 876 images from natural scenarios are collected for the test set of the MTVQA benchmark.
+To ensure the visual-textual alignment, for text-rich images lacking text and language annotations, we subject them to a standardized data cleaning process, which includes multilingual text recognition and language classification.
+Afterwards, we organize all the text-rich images we have obtained into language-specific groups, preparing them for the subsequent stage of data annotation.
 3.2 Human Expert Annotation
-In order to obtain informative and accurate text-related QA pairs on the language-specific grouped images, we recruit a group of annotators with expertise from local regions of each language.
-It is worth noting that all these annotators are native speakers of their respective languages, ensuring their deep understanding and proficiency in the linguistic nuances and cultural context necessary for precise annotations.
-Considering the subjective nature of the text-image understanding task, we have implemented a further division within the annotation team.
-This division involves separating the team into two independent groups, with one group dedicated to generating and responding to questions based on the provided images, while the other group focuses on evaluating and correcting the QA pair results.
-This raise-then-correct paradigm ensures a comprehensive and reliable assessment of the text-image understanding process.
-Additionally, each language’s annotation results undergo a 10% sampling inspection by a quality inspector.
-If the QA pairs fail to meet the criteria, they are sent back for re-annotation.
-Prior to commencing the formal human expert annotation task, all annotators undergo unified training and receive annotation examples.
-The brief diagram of the two-round annotation process is shown in Figure 3 and we elaborate on it in the following subsections.
-Table 1: Mean lengths of question-answer pairs in different languages of training set and test set, using GPT-4o tokenizer.
+To obtain informative and accurate text-related QA pairs for images grouped by specific languages, a specialized group of annotators with expertise in the local regions of each language is recruited.
+These annotators are native speakers, guaranteeing a profound understanding and skillfulness in the linguistic subtleties and cultural contexts required for precise annotations.
+Given the subjective nature of understanding text within images, the annotation team is divided into two independent groups.
+One group is tasked with generating questions and providing answers based on the images, while the other group is responsible for evaluating and correcting the QA pairs.
+This division, known as the raise-then-correct paradigm, ensures a thorough and trustworthy evaluation of the text-rich image comprehension process.
+Moreover, the annotation results for each language underwent a 10% sampling inspection by a quality inspector to ensure adherence to standards.
+QA pairs that are not meeting the criteria are returned for re-annotation.
+Before beginning the formal annotation task, all annotators have received training and examples of annotations.
+The two-round annotation process is briefly illustrated in Fig.
+4, further details of which are elaborated in the subsequent sections.
+Refer to caption Refer to caption Refer to caption
+Refer to caption Refer to caption Refer to caption
+Refer to caption Refer to caption Refer to caption
+Figure 5: Word clouds showcasing top answers in various languages, tokenized via NLTK with removing stop words, punctuation, and digits.
+First Round Questioning and Answering.
+In the first round, we allocate three annotators per language to generate initial QA results.
+The annotators are tasked with thoroughly examining the textual and visual elements within a text-rich image in our collection.
+They are to examine the textual and visual elements within the image to formulate five meaningful and distinct questions and provide corresponding answers.
+All annotators should adhere to the following criteria: (1) the first three questions should satisfy that answering these questions requires direct reading of the textual information in the image, (2) the fourth and fifth questions require reasoning about the text in the image to answer, (3) the questions and answers must be reasonably correct and consistent with the content of the image, and (4) the answer should be as concise as possible and free of nonsense (e.g., when the question is “When is the volunteer recruitment period”, the answer should be “9:00-16:00” rather than “The volunteer recruitment period is 9:00-16:00”).
+This requirement for brevity aims to make the evaluation process user-friendly and reliable, avoiding the influence of extraneous content on the evaluation metrics.
+Second Round Evaluation and Correction.
+To reduce the effect of human subjective cognitive bias on our MTVQA benchmark and get high-quality question-answer pairs, we assign two annotators for each language to carry out the annotation evaluation and correction process.
+These annotators follow a specific evaluation and correction protocol to ensure consistency and accuracy based on the first round results: (1) Assess the relevance of the question to the text in the image.
+Irrelevant QA pairs are discarded; (2) Verify the correctness of the answer and make necessary modifications; (3) Check for redundancy in the answer.
+If the answer repeats the question’s information, the repeated content is removed for conciseness.
+(4) Ethical Assessment.
+Review the image content or the question-answer pair for any unethical content, including but not limited to politics, personal privacy issues, etc.
+Such content is removed to uphold ethical standards in the dataset.
+Refer to caption Refer to caption Refer to caption
+Refer to caption Refer to caption Refer to caption
+Refer to caption Refer to caption Refer to caption
+Figure 6: Statistics of question and answer lengths of different languages aggregating training and test set, using GPT-4o tokenizer.
+3.3 Data Statistics
+The MTVQA benchmark consists of 8,794 images and 28,607 question-answer pairs across the nine languages, divided into a training set with 6,678 images and 21,829 question-answer pairs and a test set with 2,116 images and 6,778 question-answer pairs.
+Detailed data distribution is illustrated in Fig.
+2.
+Additionally, the benchmark showcases the vocabulary richness for each language through word clouds, as depicted in Fig.
+5, and the lengths of questions and answers are statistically analyzed using the GPT-4o tokenizer, as shown in Fig.
+6.
+4 Experiments
+4.1 Baseline Models
+To comprehensively assess MLLMs’ multilingual perception and comprehension capabilities, We select state-of-the-art MLLMs from three categories: open-source general MLLMs, open-source text-centric MLLMs, and closed-source MLLMs.
+Each category contains the following models: (1) General MLLMs: InternVL-V1.5 [47], InternLM-Xcomposer2-4KHD [16], Mini-Gemini-HD-34B [17], Llava-Next-34B [18], DeepSeek-VL [8], YI-VL-34B [9]; (2) Text-centric MLLMs: TextSquare [14], TextMonkey [13], mPLUG-DocOwl 1.5 [12], MiniCPM-V 2.0 [48]; (3) Closed-source MLLMs: GPT-4o [1], GPT-4V [2], Gemini Ultra [4], QwenVL Max [7], QwenVL Plus [7], Claude3 Opus [5], Claude3 Sonnet [5], and GLM-4V [49].
+For the closed-source MLLMs, we use the chat version through official APIs, while for the open-source MLLMs, we utilize the instruct versions available on the HuggingFace Model Hub.
+The open-source MLLMs’ model size varies from 7b to 34b.
+4.2 Implementation Details
+We conduct the evaluation experiments over the baseline MLLMs with their default settings, ignoring the effect of generation configuration on the results.
+To make the output of MLLMs more evaluation-friendly, we design the following prompt format to limit the output length: “Answer the question using a word or phrase in the language of the question.
++ <Question>”, where “<Question>” represents the actual question from the MTVQA test set.
+This approach aims to make the answers as concise as possible.
+Besides, InternLM-Xcomposer2-4KHD  [16] is chosen as the base model for an instruction tuning experiment on the MTVQA training set.
+The instruction tuning process adheres to the default training settings specified by the source, with “HD-16” and completes one epoch of training on 8 NVIDIA-A100 GPUs within 2 hours.
+Table 1: Performance of the leading closed- and open-source MLLMs on the MTVQA benchmark.
+The best results of each language are bolded.
+The best results for open-source models are underlined.
+“Xcomposer-SFT” denotes instruction tuning to Xcomposer2-4KHD with MTVQA’s training set.
+AR DE FR IT JA KO RU TH VI Avg.
+Closed-Source MLLMs
+GPT-4V [2] 11.5 31.5 40.4 32.3 11.5 16.7 10.3 15.0 28.9 22.0
+GPT-4o [1] 20.2 34.2 41.2 32.7 20.0 33.9 11.5 22.5 34.2 27.8
+Gemini Ultra [4] 14.7 32.3 40.0 31.8 12.3 17.2 11.8 20.3 28.6 23.2
+QwenVL Max [7] 7.7 31.4 37.6 30.2 18.6 25.4 10.4 4.8 23.5 21.1
+QwenVL Plus [7] 4.8 28.8 33.7 27.1 12.8 19.9 9.4 5.6 18.1 17.8
+Claude3 Opus [5] 15.1 33.4 40.6 34.4 19.4 27.2 13.0 19.5 29.1 25.7
+Claude3 Sonnet [5] 10.5 28.9 35.6 31.8 13.9 22.2 11.0 15.2 20.8 21.1
+GLM-4V [49] 0.3 30.0 34.1 30.1 3.4 5.7 3.0 3.5 12.3 13.6
+Open-Source MLLMs
+InternVL-V1.5 [47] 3.4 27.1 31.4 27.1 9.9 9.0 4.9 8.7 12.4 14.9
+Mini-Gemini-HD-34B [17] 2.2 25.0 29.2 25.5 6.1 8.6 4.1 4.3 11.8 13.0
+Llava-Next-34B [18] 3.3 24.0 28.0 22.3 3.6 6.1 2.6 0.4 9.8 11.1
+DeepSeek-VL [8] 0.6 14.2 15.3 15.2 2.9 3.8 1.6 0.9 5.2 6.6
+YI-VL-34B [9] 1.7 13.5 15.7 12.1 4.8 5.2 0.8 3.5 4.1 6.8
+MiniCPM-V 2.0 [48] 1.3 12.7 14.9 17.0 3.7 5.6 2.2 2.2 6.8 7.4
+MiniCPM-V 2.5 [48] 6.1 29.6 35.7 26.0 12.1 13.1 5.3 12.6 15.3 17.3
+TextSquare [14] 3.7 27.0 30.8 26.7 3.2 7.2 6.7 5.2 12.4 13.6
+TextMonkey [13] 2.0 18.1 19.9 22.1 4.6 7.2 3.2 0.9 11.1 9.9
+mPLUG-DocOwl 1.5 [12] 1.0 13.9 14.9 18.2 2.9 5.0 2.0 0.9 6.4 7.2
+Xcomposer2-4KHD [16] 2.0 20.6 23.2 21.6 5.6 7.7 4.1 6.1 10.1 11.2
+Xcomposer-SFT 11.8 31.7 37.4 29.3 14.5 12.9 5.8 13.9 20.2 19.7
+4.3 Evaluation Results
+Evaluation metric.
+To accurately assess whether the visual text that occurs in the answer is correct, we adopt Accuracy as the metric.
+The Accuracy metric measures the percentage of questions for which the predicted answer matches exactly with any of the target answers for the question.
+The accuracy metric awards a zero score even when the prediction differs slightly from the target answer.
+[…]
+[…]
+Table 2: Few-shot performance of GPT-4V on the MTVQA benchmark.
+In-context examples are randomly selected from the training set of MTVQA in the respective languages.
+“n-shot” represents the number of the selected in-context examples.
+AR DE FR IT JA KO RU TH VI Avg.
+zero-shot 11.5 31.5 40.4 32.3 11.5 16.7 10.3 15.0 28.9 22.0
+two-shot 11.6 35.3 42.2 33.1 13.2 17.0 11.4 19.0 31.0 23.7
+five-shot 13.5 34.8 43.0 35.7 13.4 18.8 11.6 19.2 33.0 24.8
+eight-shot 15.5 34.8 42.1 35.6 13.4 17.5 12.1 18.2 32.8 24.7
+Few-shot evaluation of GPT-4V.
+Here, we compare the performance of GPT-4V on MTVQA under the few-shot settings.
+Specifically, we perform zero-shot, two-shot, five-shot, and eight-shot evaluations.
+We randomly select in-context examples from the train set in the respective languages and evaluate GPT-4V on the remaining instances.
+As shown in Tab.
+2, compared to the zero-shot setting, GPT-4V’s performance has improved considerably under few-shot, highlighting its exceptional in-context learning ability in multilingual text comprehension contexts.
+Moreover, comparisons based on varying numbers of in-context samples reveal that augmenting the in-context samples can aid in further enhancement; however, after reaching a certain volume, the improvement becomes saturated.
+5 Limitation
+The current version of the MTVQA dataset, while dialectically diverse, exhibits limitations in its language coverage.
+Despite encompassing a range of languages, it falls short of inclusivity, omitting numerous lesser-spoken languages.
+This prompts our future continual endeavor to ensure comprehensive representation across the linguistic spectrum.
+Additionally, the dataset currently provides a single canonical response per question, which may not fully capture the range of answers corresponding to different expressions of the same underlying semantics.
+Recognizing the multifaceted nature of the inquiries, future versions will aim to include a spectrum of plausible answers to reflect the varied perspectives inherent to each question.
+6 Conclusion
+To attack the visual-textual misalignment issue in multilingual TEC-VQA, we introduce MTVQA, a new benchmark featuring high-quality human expert annotations in 9 diverse languages.
+We believe that MTVQA is the first benchmark to provide fully manual annotations tailored to text-centric scenarios.
+The results obtained from closed-source, general-purpose and text-centric MLLMs on our MTVQA dataset indicate that there is still room for improving their performance in multilingual text-centric scenarios.
+Although the MTVQA dataset has limitations, such as the underrepresentation of several lesser-spoken languages and single canonical answers, future updates will address these issues by expanding the multilingual scope and including a range of plausible answers.
+We are confident that this dataset can inspire researchers within the TEC-VQA community with new perspectives and ideas.
+Appendix A Appendix
+(a) German (DE) example
+(b) Italian (IT) example
+(c) French (FR) example
+(d) Japanese (JA) example
+(e) Korean (KO) example
+Figure 7: Multilingual VQA examples selected from five languages. The corresponding translations in English are in brackets.
+Table 3: Mean lengths of question-answer pairs in different languages of the training set and test set, using GPT-4o tokenizer.
 AR DE FR IT JA KO RU TH VI
@@
 Answer 7.95 6.67 6.61 11.04 12.55 13.61 14.42 12.08 13.00
-First Round Questioning and Answering.
-For the first round of annotation tasks, we assigned 3 annotators for each language to manually generate original QA results.
-Given a text-centric image from our collection, annotators are first required to read the texts in the image and analyze other contents in the image in a comprehensive and detailed manner.
-They must then raise 4 meaningful and distinct questions based on the content in the image and give the answers.
-All annotators adhere to the following criteria: (1) the first three questions should satisfy that answering these questions requires direct reading of the textual information in the image, (2) the fourth question requires reasoning about the text in the image to answer (3) the questions and answers must be reasonably correct and consistent with the content of the image, and (4) the answer should be as concise as possible and free of nonsense (e.g., when the question is “When is the volunteer recruitment period”, the answer should be “9:00-16:00” rather than “The volunteer recruitment period is 9:00-16:00”).
-It’s worth mentioning that our requirement for concise answers is to make the evaluation process more friendly and more reliable, cause we try to keep the evaluation metrics unaffected by extraneous content in the answer sentence.
-Second round Evaluation and Correction.
-To reduce the effect of human subjective cognitive bias on our MTVQA benchmark and get high-quality question-answer pairs, we assigned 2 annotators for each language for the annotation evaluation and correction process.
-Based on the provided images and the first-round annotation results, the annotators must follow these rules of judgment and steps for the annotation: (1) Whether the question is related to the text in the image.
-If not, discard the current question-answer pair, (2) Whether the answer is correct.
-If not, modify the answer, and (3) Whether the answer repeats the content from the question.
-If so, remove the repeated content to ensure a concise answer.
-Figure 3: A brief diagram of the annotation process.
-3.3 Data Statistics
-We instruct the annotators to complete the above human expert annotation work towards the text-centric VQA tasks and construct the MTVQA benchmark consisting of 8,794 images and 28,607 question-answer pairs that cover the 9 languages.
-The MTVQA benchmark is divided into a training set containing 6,678 images and 21,829 question-answer pairs, and a test set containing 2,116 images and 6,778 question-answer pairs.
-The detailed data distribution can be seen in Figure 1.
-To visualize the vocabulary richness of our benchmark, we calculate the word frequencies for each language and present them in the form of word clouds as shown in Figure 4.
-In Figure 5 we demonstrate the statistics of the question and answer lengths using GPT-4o tokenizer.
 Refer to caption Refer to caption Refer to caption
@@
 Refer to caption Refer to caption Refer to caption
-Figure 4: Top answer word clouds of different languages.
-Refer to caption Refer to caption Refer to caption
-Refer to caption Refer to caption Refer to caption
-Refer to caption Refer to caption Refer to caption
-Figure 5: Statistics of question and answer lengths of different languages aggregating training and test set, using GPT-4o tokenizer.
-4 Experiments
-4.1 Baseline Models
-For the MTVQA benchmark, we evaluate the following instruction-tuned general MLLMs, (1) Open-source MLLMs: InternVL-V1.5 [46], InternLM-Xcomposer2-4KHD [15], Mini-Gemini-HD-34B [16], Llava-Next-34B [17], DeepSeek-VL [7], YI-VL-34B [8], TextSquare [13], TextMonkey [12] and mPLUG-DocOwl 1.5 [11]; (2) Closed-source MLLMs: GPT-4V, Gemini Ultra, QwenVL Max, QwenVL Plus, Claude3 Opus, Claude3 Sonnet and GLM4V.
-For the closed-source MLLMs, we use the chat version through the official APIs, while for the open-source MLLMs, we utilize the instruct versions.
-It is noted that all the model weights of the open-source MLLMs evaluated in our experiments could be downloaded from the HuggingFace Model Hub.
-For the open-source MLLMs, the model size varies from 7b to 34b.
-4.2 Implementation Details
-We conduct the evaluation experiments over the baseline MLLMs with their default settings, ignoring the effect of generation configuration on the results.
-To make the output of MLLMs more evaluation-friendly, we design the following prompt format to limit the output length: “Answer the question using a word or phrase in the language of the question.
-+ <Question>”, where <Question> represents the corresponding question of the input image.
-The extra prefixes added to the raw question could limit the answer to be as concise as possible.
-Besides, we utilize the InternLM-Xcomposer2-4KHD [15] as the backbone for the instructional fine-tuning experiment on the MTVQA training set.
-In the instructional fine-tuning process, we follow the default training settings [15] with “HD-16” and train on MTVQA training set for 1 epoch.
-4.3 Evaluation Results
-[…]
-[…]
-Table 2: Performance of the leading closed-source and open-source MLLMs on the MTVQA benchmark.
-AR DE FR IT JA KO RU TH VI Avg.
-Closed-Source MLLMs
-GPT-4V 11.5 31.5 40.4 32.3 11.5 16.7 10.3 15.0 28.9 22.0
-Gemini Ultra 14.7 32.3 40.0 31.8 12.3 17.2 11.8 20.3 28.6 23.2
-QwenVL Max 7.7 31.4 37.6 30.2 18.6 25.4 10.4 4.8 23.5 21.1
-QwenVL Plus 4.8 28.8 33.7 27.1 12.8 19.9 9.4 5.6 18.1 17.8
-Claude3 Opus 15.1 33.4 40.6 34.4 19.4 27.2 13.0 19.5 29.1 25.7
-Claude3 Sonnet 10.5 28.9 35.6 31.8 13.9 22.2 11.0 15.2 20.8 21.1
-GLM4V 0.3 30.0 34.1 30.1 3.4 5.7 3.0 3.5 12.3 13.6
-Open-Source MLLMs
-InternVL-V1.5 [46] 3.4 27.1 31.4 27.1 9.9 9.0 4.9 8.7 12.4 14.9
-Mini-Gemini-HD-34B [16] 2.2 25.0 29.2 25.5 6.1 8.6 4.1 4.3 11.8 13.0
-Llava-Next-34B [17] 3.3 24.0 28.0 22.3 3.6 6.1 2.6 0.4 9.8 11.1
-DeepSeek-VL [7] 0.6 14.2 15.3 15.2 2.9 3.8 1.6 0.9 5.2 6.6
-YI-VL-34B [8] 1.7 13.5 15.7 12.1 4.8 5.2 0.8 3.5 4.1 6.8
-TextSquare [13] 3.7 27.0 30.8 26.7 3.2 7.2 6.7 5.2 12.4 13.6
-TextMonkey [12] 2.0 18.1 19.9 22.1 4.6 7.2 3.2 0.9 11.1 9.9
-mPLUG-DocOwl 1.5 [11] 1.0 13.9 14.9 18.2 2.9 5.0 2.0 0.9 6.4 7.2
-Xcomposer2-4KHD [15] 2.0 20.6 23.2 21.6 5.6 7.7 4.1 6.1 10.1 11.2
-Xcomposer-SFT 11.8 31.7 37.4 29.3 14.5 12.9 5.8 13.9 20.2 19.7
-5 Limitation
-The current iteration of MTVQA exhibits certain constraints that warrant attention.
-Primarily, the linguistic diversity incorporated is not exhaustive; several lesser-spoken languages remain unrepresented.
-Future enhancements will aim to broaden the multilingual scope of the dataset.
-Additionally, the dataset currently offers a singular canonical response for each question.
-Recognizing the multifaceted nature of the inquiry, subsequent versions will endeavor to include a spectrum of plausible answers to reflect the varied perspectives inherent to each question.
-6 Conclusion
-In this paper, we introduce MTVQA, a multilingual TEC-VQA benchmark featuring high-quality human expert annotations in 9 diverse languages.
-We believe that MTVQA is the first benchmark of its kind to provide fully manual annotations specifically tailored to text-centric scenarios.
-The results obtained from both closed- and open-source MLLMs on our MTVQA dataset indicate that there is still room for improving their performance in multilingual text-centric scenarios.
-Although the current version of MTVQA has constraints regarding linguistic diversity and singular responses per question, we are confident that this dataset can still inspire researchers within the TEC-VQA community with new perspectives and ideas.
+Figure 8: Word clouds showcasing top questions in various languages, tokenized via NLTK with the removal of stop words, punctuation, and digits.

### v3

@@
 Text-Centric Visual Question Answering (TEC-VQA) in its proper format not only facilitates human-machine interaction in text-centric visual environments but also serves as a de facto gold proxy to evaluate AI models in the domain of text-centric scene understanding.
-Nonetheless, most existing TEC-VQA benchmarks have focused on high-resource languages like English and Chinese.
-Despite pioneering works to expand multilingual QA pairs in non-text-centric VQA datasets through translation engines, the translation-based protocol encounters a substantial “visual-textual misalignment” problem when applied to TEC-VQA.
+Nonetheless, most existing TEC-VQA benchmarks focus on high-resource languages like English and Chinese.
+Despite pioneering works expanding multilingual QA pairs in non-text-centric VQA datasets through translation engines, the translation-based protocol encounters a substantial “visual-textual misalignment” problem when applied to TEC-VQA.
 Specifically, it prioritizes the text in question-answer pairs while disregarding the visual text present in images.
@@
 In this work, we tackle multilingual TEC-VQA by introducing MTVQA, the first benchmark featuring high-quality human expert annotations across 9 diverse languages, consisting of 6,778 question-answer pairs across 2,116 images.
-Further, by comprehensively evaluating numerous state-of-the-art Multimodal Large Language Models (MLLMs), including GPT-4o, GPT-4V, Claude3, and Gemini, on the MTVQA dataset, it is evident that there is still a large room for performance improvement, underscoring the value of MTVQA.
+Further, by comprehensively evaluating numerous state-of-the-art Multimodal Large Language Models (MLLMs), including Qwen2-VL, GPT-4o, GPT-4V, Claude3, and Gemini, on the MTVQA benchmark, it is evident that there is still a large room for performance improvement (Qwen2-VL scoring 30.9 versus 79.7 for human performance), underscoring the value of MTVQA.
 Additionally, we supply multilingual training data within the MTVQA dataset, demonstrating that straightforward fine-tuning with this data can substantially enhance multilingual TEC-VQA performance.
 We aspire that MTVQA will offer the research community fresh insights and stimulate further exploration in multilingual visual text comprehension.
-The project homepage is at https://bytedance.github.io/MTVQA/.
-footnotetext: Equal contribution. 🖂 Corresponding authors.
 1 Introduction
 In the era of burgeoning AI, especially in MLLMs [34, 1, 47, 44, 3, 39, 5, 27, 49, 12, 13, 18, 26, 43, 9, 11, 22, 24], Text-Centric Visual Question Answering (TEC-VQA[6, 42, 13, 12, 43, 26, 18] has served as a de facto gold proxy to evaluate AI models in the domain of text-centric scene understanding.
-Compared with general VQA [19, 21, 22, 20, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34], TEC-VQA places greater emphasis on answering questions that require understanding textual information within images.
-It provides a streamlined avenue for individuals without specialized expertise to articulate their requirements and access applications in text-centric visual environments.
-However, the majority of advancements in TEC-VQA have predominantly concentrated on high-resource languages, e.g., English [19, 20, 21, 24], Chinese [35, 32], Japanese [36, 37] and etc., thus restricting the applicability of AI models to the global community, particularly populations speaking low-resource languages.
-To tackle the problem of language diversity, several seminal studies [38, 39, 40] in the general VQA field, leverage off-the-shelf translation engines to expand existing question-answer pairs from high-resource languages to their multilingual counterparts including low-resource ones.
-However, when applied to TEC-VQA, this translation-based approach may fall prey to the “visual-textual misalignment” problem as only the text in question-answer pairs can be processed, while the visual text present in the images is overlooked.
-Although the visual text can be obtained by an OCR (Optical Character Recognition) engine and then translated into English, this indirect process could introduce issues such as nuanced meaning, contextual distortion, language bias, and question type diversity.
-These factors further render the transferability of the translation protocol infeasible for TEC-VQA.
+Compared with general VQA [6, 30, 36, 42, 32, 31, 29, 52, 20, 4, 28, 40, 25, 15, 14, 23], TEC-VQA places greater emphasis on answering questions that require understanding visual textual information within images.
+It enables individuals without specialized expertise to access applications in text-centric visual environments.
+However, most advancements in TEC-VQA have predominantly concentrated on high-resource languages, e.g., English [6, 42, 30, 31], Chinese [37, 15], Japanese [41, 33] and etc., thus restricting the applicability of AI models to the global community, particularly populations speaking low-resource languages.
+To tackle the problem of language diversity, several seminal studies [38, 35, 7] in the general VQA field simply leverage translation engines to expand question-answer pairs from high-resource to low-resource languages.
+However, this translation-based approach is not feasible for TEC-VQA, as it merely processes text in question-answer pairs, neglecting the critical visual text needed for answering based on images.
+Although the visual text can be recognized by an OCR (Optical Character Recognition) engine and then translated into the target language, this indirect process could lead to a “visual-textual misalignment” problem, due to nuanced meaning, contextual distortion, language bias, and question type diversity.
 Taking the second case in Fig.
-1 as an example, if either the visual text or the question in Russian is recognized or translated problematically, then the question would never be answered.
-The status quo begs for a question: “Can we address the visual-textual misalignment problem directly from the source language per se for multilingual TEC-VQA and what we stand in the MLLM era?
+1 as an example, if either the visual text or the question in Russian is recognized or translated problematically, then the question would never be answered correctly.
+The status quo begs for a question: “Can we directly leverage visual text in source language per se for multilingual TEC-VQA and what we stand in the MLLM era?
 Figure 1: Multilingual text-centric VQA visualization selected from four languages.
@@
 The corresponding translations in English are in brackets.
-More examples see Figure 7.
 Refer to caption Refer to caption
@@
 We further investigate recent representative MLLMs in Fig.
-3, including GPT-4o, GPT-4V, Gemini, QwenVL, etc., and their performance on our proposed MTVQA.
+3, including GPT-4o, GPT-4V, Gemini, Qwen2-VL, etc., and their performance on MTVQA.
 As vividly revealed in Fig.
-3 (a), GPT-4o is the top performer, followed by Claude3 Opus and Gemini Ultra, with closed-source MLLMs generally leading.
-General MLLMs like QwenVL Max and QwenVL Plus show mid-tier performance, while text-centric MLLMs lag behind.
+3 (a), the general MLLM Qwen2-VL is the top performer, followed by closed-source GPT-4o and Claude3 Opus.
+Other general MLLMs like QwenVL Max and QwenVL Plus show mid-tier performance, while text-centric MLLMs lag behind.
 Fig.
@@
 The results unequivocally demonstrate that opportunities for improvement persist within existing MLLMs when applied in multilingual text-rich scenarios.
-In summary, the main contributions of this paper can be categorized into three points:
-  • We coin the MTVQA dataset, which to the best of our knowledge is the first multilingual TEC-VQA benchmark providing human expert annotations to solve “visual-textual misalignment” problem in multilingual text-centric scenarios.
-  • We benchmark the state-of-the-art MLLMs on our new dataset and show that there are still opportunities for improvement on even the most advanced MLLMs when it comes to multilingual text-rich scenarios.
-  • We establish a set of new multilingual TEC-VQA baselines for both closed-source and general-purpose, text-centric MLLMs.
+In summary, the main contributions of this paper can be categorized into four points:
+  • We coin the MTVQA dataset, which, to the best of our knowledge, is the first multilingual TEC-VQA benchmark providing human expert annotations to solve the “visual-textual misalignment” problem in multilingual text-centric scenarios.
+  • We benchmark various MLLMs and human performance on MTVQA and show that there are still great opportunities for improvement on even the most advanced MLLMs in multilingual text-rich scenarios.
+  • We establish a set of new multilingual TEC-VQA baselines for closed-source and general-purpose, text-centric MLLMs.
+  • MTVQA has been serving as a standard benchmark for evaluating multilingual visual text comprehension capabilities by some recently released state-of-the-art MLLMs such as Qwen2-VL [46] and InternVL 2 [11].
 2 Related Work
+Table 1: Comprehensive comparison on the benchmarks related to MTVQA. GPT4V denotes GPT-4V performance.
+Benchmark Scene Manual QA QA Language Visual Text Language GPT4V
+OCRBench Multiple Text-rich English English 64.5%
+TextVQA Scene Text English English 78.0%
+DocVQA Document English English 88.4%
+EST-VQA Scene Text English, Chinese English, Chinese 72.3%
+xGQA General 7 Languages - 67.7%
+MaXM General 7 Languages - 62.8%
+MTVQA Multiple Text-rich 9 Languages 10 languages 22.0%
 2.1 MLLMs for Text-centric VQA
-Recent advancements in MLLMs [2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18] have revolutionized VQA tasks, as demonstrated by the remarkable zero-shot performance of these models.
+Recent advancements in MLLMs [1, 47, 44, 3, 39, 5, 27, 49, 12, 13, 18, 26, 43, 9, 11, 22, 24, 51] have revolutionized VQA tasks, as demonstrated by the remarkable zero-shot performance of these models.
 Notably, the high generalizability of MLLMs, when explicitly trained on visual text understanding datasets and fine-tuned with instructions, has significantly enhanced their application in text-centric VQA scenarios [13, 12, 43, 26, 18].
@@
 One notable limitation of these datasets is their predominant focus on English [6, 42, 30, 31] or other high-resource languages such as Chinese [37, 15] and Japanese [41, 33], which restricts the applicability of VQA systems for low-resource languages such as Thai and Vietnamese.
-There are some recent efforts toward extending VQA tasks to a wider range of languages [43, 39, 44, 40, 45, 38] by providing a multilingual VQA datasets.
+There are some recent efforts toward extending VQA tasks to a broader range of languages [17, 35, 45, 7, 21, 38] by providing a multilingual VQA datasets.
 For example,  Gao et al.
@@
 [38] developed a large-scale multilingual and code-mixed VQA dataset (MuCo-VQA) supporting five languages.
-Of more relevance are the works xGQA (8 languages) [39] and MaXM (7 languages) [40], which apply translation-based protocols to expand VQA data beyond English.
+Of more relevance are the works xGQA (7 languages) [35] and MaXM (7 languages) [7], which apply translation-based protocols to expand VQA data beyond English.
 However, the translation-based multilingual VQA datasets inherently face issues, such as the “visual-textual misalignment” problem, where only the textual information in question-answer pairs is considered, while the visual text in images is overlooked.
@@
 Our benchmark, MTVQA, distinguishes itself by focusing on multilingual text-centric VQA scenarios using human expert annotations.
-It covers 9 languages, thereby facilitating the training and evaluation of multilingual models in diverse linguistic contexts.
+It covers 9 languages, facilitating the training and evaluation of multilingual models in diverse linguistic contexts.
 Additionally, our dataset can gauge the VQA system’s ability for not only high-resource languages but also those that are typically underrepresented in current datasets [6, 42, 30, 31, 15].
@@
 MTVQA covers nine languages: Arabic (AR), Korean (KO), Japanese (JA), Thai (TH), Vietnamese (VI), Russian (RU), French (FR), German (DE), and Italian (IT).
-The data construction process involves a comprehensive collection of text-rich images and a two-round human expert annotation process.
-Figure 4: A brief diagram of the annotation process.
+The data construction involves a comprehensive collection of text-rich images and a two-round human expert annotation process.
+In terms of data volume, an initial dataset of 8,980 images is compiled.
+Following a data cleaning phase, 8,895 images are retained for annotation.
+The first round of annotation yielded 8,895 images, corresponding to 28,906 question-answer (QA) pairs.
+After the second round of annotation and final quality control measures, the final dataset comprises 8,794 images and 28,607 QA pairs.
+The benchmark construction cost is divided into two parts: image acquisition and annotation.
+Image acquisition costs about two months and 30,000 dollars.
+VQA annotation costs about three months and 60,000 dollars.
+Hourly wages vary from country to country, with senior language specialists costing between 20 and 40 per hour and local annotators costing between 8 and 20 per hour.
+The average time spent by each labeler was roughly 60 days.
+Figure 4: A brief pipeline of the annotation process.
 3.1 Data Collection
-The raw data collection is oriented towards gathering text-rich images from various natural and document scenarios, ensuring diversity and quality.
-This includes images from publicly available datasets(e.g., ICDAR MLT19 [46]) and those sourced from the internet, such as menus, logos, maps, bills, PowerPoint slides (PPTs), research papers, etc, as shown in Fig.
+The raw data collection aims to gather text-rich images from various scene text and document scenarios, ensuring diversity and quality.
+This includes images from publicly available datasets (e.g., ICDAR MLT19 [16]) and those sourced from the internet (i.e., Laion-OCR, which is filtered from Common Crawl [10]), such as menus, logos, maps, bills, PowerPoint slides (PPTs), research papers, etc, as shown in Fig.
 2.
+The image collection process consists of two steps: extracting multilingual text using a multilingual OCR engine (Bytedance multilingual OCR API [^1] https://cv-api.bytedance.com/api/ocr-process/v1/multi_language_v2) and selecting by language types and the amount of text contained in the image.
+Approximately 30% of the overall image data is obtained from public datasets, with 20% sourced from the web and 50% from manual collection, respectively.
+Manual collection means that the images are taken in different scenarios in the countries and areas where the corresponding language is spoken, which means that the images are native and high quality.
 A total of 1,220 images from document scenarios and 876 images from natural scenarios are collected for the test set of the MTVQA benchmark.
-To ensure the visual-textual alignment, for text-rich images lacking text and language annotations, we subject them to a standardized data cleaning process, which includes multilingual text recognition and language classification.
-Afterwards, we organize all the text-rich images we have obtained into language-specific groups, preparing them for the subsequent stage of data annotation.
+To ensure data content meets regulatory requirements, we subject them to a standardized data cleaning process.
+The image-cleaning processing pipeline involves a preliminary round of algorithm-driven review to identify and filter out unusable images.
+It includes detecting and removing images with politically sensitive, pornographic, violent, or other undesirable features.
+Subsequently, the multilingual OCR tool (Bytedance multilingual OCR API) is employed to extract the textual content from the remaining images.
+Images devoid of textual information are discarded, and the surviving images are categorized based on their language.
+Afterward, we organize all the text-rich images we have obtained into language-specific groups, preparing them for the subsequent stage of data annotation.
 3.2 Human Expert Annotation
 To obtain informative and accurate text-related QA pairs for images grouped by specific languages, a specialized group of annotators with expertise in the local regions of each language is recruited.
-These annotators are native speakers, guaranteeing a profound understanding and skillfulness in the linguistic subtleties and cultural contexts required for precise annotations.
+The annotators must be native speakers of the corresponding language and have actively utilized it for a minimum duration of 10 years.
+Additionally, they must possess at least a university-level degree or higher academic qualification, guaranteeing a profound understanding and skillfulness in the linguistic subtleties and cultural contexts required for precise annotations.
 Given the subjective nature of understanding text within images, the annotation team is divided into two independent groups.
@@
 Moreover, the annotation results for each language underwent a 10% sampling inspection by a quality inspector to ensure adherence to standards.
-QA pairs that are not meeting the criteria are returned for re-annotation.
-Before beginning the formal annotation task, all annotators have received training and examples of annotations.
+QA pairs that do not meet the criteria are returned for re-annotation.
+Prior to the formal annotation process, all annotators will be provided with a detailed explanation of the annotation guidelines, including a dedicated question-and-answer session to clarify any ambiguities and uncertainties.
+A pilot annotation task will be conducted on a limited dataset to ensure a shared understanding of the annotation rules.
 The two-round annotation process is briefly illustrated in Fig.
 4, further details of which are elaborated in the subsequent sections.
-Refer to caption Refer to caption Refer to caption
-Refer to caption Refer to caption Refer to caption
-Refer to caption Refer to caption Refer to caption
-Figure 5: Word clouds showcasing top answers in various languages, tokenized via NLTK with removing stop words, punctuation, and digits.
 First Round Questioning and Answering.
@@
 Second Round Evaluation and Correction.
-To reduce the effect of human subjective cognitive bias on our MTVQA benchmark and get high-quality question-answer pairs, we assign two annotators for each language to carry out the annotation evaluation and correction process.
-These annotators follow a specific evaluation and correction protocol to ensure consistency and accuracy based on the first round results: (1) Assess the relevance of the question to the text in the image.
+To reduce the effect of human subjective cognitive bias on our MTVQA benchmark and get high-quality question-answer pairs, we assign two annotators for each language to carry out the annotation evaluation and correction process independently.
+These annotators follow a specific evaluation and correction protocol to ensure consistency and accuracy based on the first round results: (1) Assess the question’s relevance to the text in the image.
 Irrelevant QA pairs are discarded; (2) Verify the correctness of the answer and make necessary modifications; (3) Check for redundancy in the answer.
@@
 Figure 5: Statistics of question and answer lengths of different languages aggregating training and test set, using GPT-4o tokenizer.
+Refer to caption Refer to caption Refer to caption
+Refer to caption Refer to caption Refer to caption
+Refer to caption Refer to caption Refer to caption
+Figure 6: Word clouds showcasing top questions in various languages, tokenized via NLTK with the removal of stop words, punctuation, and digits.
 3.3 Data Statistics
@@
 Additionally, the benchmark showcases the vocabulary richness for each language through word clouds, as depicted in Fig.
-5, and the lengths of questions and answers are statistically analyzed using the GPT-4o tokenizer, as shown in Fig.
-6.
+8, and the lengths of questions and answers are statistically analyzed using the GPT-4o tokenizer, as shown in Fig.
+5.
 4 Experiments
@@
 To comprehensively assess MLLMs’ multilingual perception and comprehension capabilities, We select state-of-the-art MLLMs from three categories: open-source general MLLMs, open-source text-centric MLLMs, and closed-source MLLMs.
-Each category contains the following models: (1) General MLLMs: InternVL-V1.5 [47], InternLM-Xcomposer2-4KHD [16], Mini-Gemini-HD-34B [17], Llava-Next-34B [18], DeepSeek-VL [8], YI-VL-34B [9]; (2) Text-centric MLLMs: TextSquare [14], TextMonkey [13], mPLUG-DocOwl 1.5 [12], MiniCPM-V 2.0 [48]; (3) Closed-source MLLMs: GPT-4o [1], GPT-4V [2], Gemini Ultra [4], QwenVL Max [7], QwenVL Plus [7], Claude3 Opus [5], Claude3 Sonnet [5], and GLM-4V [49].
+Each category contains the following models: (1) General MLLMs: Qwen2-VL  [46], InternVL-V1.5 [8], InternLM-Xcomposer2-4KHD [11], Mini-Gemini-HD-34B [22], Llava-Next-34B [24], DeepSeek-VL [27], YI-VL-34B [49]; (2) Text-centric MLLMs: TextSquare [43], TextMonkey [26], mPLUG-DocOwl 1.5 [18], MiniCPM-V 2.0 [19]; (3) Closed-source MLLMs: GPT-4o [34], GPT-4V [1], Gemini Ultra [44], QwenVL Max [5], QwenVL Plus [5], Claude3 Opus [3], Claude3 Sonnet [3], and GLM-4V [2].
 For the closed-source MLLMs, we use the chat version through official APIs, while for the open-source MLLMs, we utilize the instruct versions available on the HuggingFace Model Hub.
-The open-source MLLMs’ model size varies from 7b to 34b.
-4.2 Implementation Details
-We conduct the evaluation experiments over the baseline MLLMs with their default settings, ignoring the effect of generation configuration on the results.
-To make the output of MLLMs more evaluation-friendly, we design the following prompt format to limit the output length: “Answer the question using a word or phrase in the language of the question.
-+ <Question>”, where “<Question>” represents the actual question from the MTVQA test set.
-This approach aims to make the answers as concise as possible.
-Besides, InternLM-Xcomposer2-4KHD  [16] is chosen as the base model for an instruction tuning experiment on the MTVQA training set.
-The instruction tuning process adheres to the default training settings specified by the source, with “HD-16” and completes one epoch of training on 8 NVIDIA-A100 GPUs within 2 hours.
+The open-source MLLMs’ model size varies from 7b to 76b.
+models GPT-4V GPT-4o Gemini Ultra Claude3 Opus Qwen2-VL 92B InternVL2 76B
+ANLS 29.6 36.9 31.7 35.4 38.7 29.1
 Table 2: Performance of the leading closed- and open-source MLLMs on the MTVQA benchmark.
 The best results of each language are bolded.
-The best results for open-source models are underlined.
+The second best results are underlined.
 “Xcomposer-SFT” denotes instruction tuning to Xcomposer2-4KHD with MTVQA’s training set.
 AR DE FR IT JA KO RU TH VI Avg.
+Human Performance
+Human 76.9 80.2 84.1 78.0 79.1 81.7 76.3 78.4 82.8 79.7
 Closed-Source MLLMs
@@
 GPT-4o [34] 20.2 34.2 41.2 32.7 20.0 33.9 11.5 22.5 34.2 27.8
+GPT-4o mini [34] 16.9 33.0 41.2 32.1 18.5 27.4 11.5 19.9 29.1 25.5
 Gemini Ultra [44] 14.7 32.3 40.0 31.8 12.3 17.2 11.8 20.3 28.6 23.2
@@
 Open-Source MLLMs
+Qwen2-VL 72B [46] 20.7 36.5 44.1 42.8 21.6 37.4 15.6 17.7 41.6 30.9
+InternVL2 76B [1] 9.5 31.3 35.7 35.2 11.1 14.3 11.9 10.0 26.9 22.0
 InternVL-V1.5 [8] 3.4 27.1 31.4 27.1 9.9 9.0 4.9 8.7 12.4 14.9
@@
 Xcomposer-SFT 11.8 31.7 37.4 29.3 14.5 12.9 5.8 13.9 20.2 19.7
-4.3 Evaluation Results
-Evaluation metric.
-To accurately assess whether the visual text that occurs in the answer is correct, we adopt Accuracy as the metric.
-The Accuracy metric measures the percentage of questions for which the predicted answer matches exactly with any of the target answers for the question.
-The accuracy metric awards a zero score even when the prediction differs slightly from the target answer.
-[…]
-[…]
+4.2 Implementation Details
+We conduct the evaluation experiments over the baseline MLLMs with their default settings, ignoring the effect of generation configuration on the results.
+To make the output of MLLMs more evaluation-friendly, we design the following prompt format to limit the output length: “Answer the question using a word or phrase in the language of the question.
++ <Question>”, where “<Question>” represents the actual question from the MTVQA test set.
+This approach aims to make the answers as concise as possible.
+Besides, InternLM-Xcomposer2-4KHD  [11] is chosen as the base model for an instruction tuning experiment on the MTVQA training set.
+The instruction tuning process adheres to the default training settings specified by the source, with “HD-16” and completes one epoch of training on 8 NVIDIA-A100 GPUs within 2 hours.
 Table 3: Few-shot performance of GPT-4V on the MTVQA benchmark.
@@
 eight-shot 15.5 34.8 42.1 35.6 13.4 17.5 12.1 18.2 32.8 24.7
+Table 4: Comparison on applying OCR to GPT-4 and GPT-4V
+AR DE FR IT JA KO RU TH VI Avg.
+OCR+GPT-4 18.9 21.2 29.8 27.3 15.8 24 10.6 19 28.2 21.6
+OCR+GPT-4V 22.3 35.1 42.6 36.7 19.2 32.1 12.1 20.9 34.1 28.3
+GPT-4V 11.5 31.5 40.4 32.3 11.5 16.7 10.3 15 28.9 22.0
+4.3 Evaluation Results
+Evaluation metric.
+To accurately assess whether the visual text that occurs in the answer is correct, we adopt Accuracy as the metric.
+The Accuracy metric measures the percentage of questions for which the predicted answer matches with any of the target answers for the question.
+We also compare the models and human performance using the ANLS and Accuracy metrics.
+The models perform a little higher (5% - 10%) when using ANLS, but the rankings are mostly unchanged.
+Since the ANLS metrics do not accurately reflect whether the correct textual content in the image is being presented, we chose the Accuracy metric.
+Human performance.
+To establish a benchmark for human performance on MTVQA, 10 well-educated native speakers are randomly selected for each language.
+As shown in Table 2, the average human performance on MTVQA is around 80%.
+[…]
+[…]
 Few-shot evaluation of GPT-4V.
@@
 As shown in Tab.
-2, compared to the zero-shot setting, GPT-4V’s performance has improved considerably under few-shot, highlighting its exceptional in-context learning ability in multilingual text comprehension contexts.
+3, compared to the zero-shot setting, GPT-4V’s performance has improved considerably under few-shot, highlighting its exceptional in-context learning ability in multilingual text comprehension contexts.
 Moreover, comparisons based on varying numbers of in-context samples reveal that augmenting the in-context samples can aid in further enhancement; however, after reaching a certain volume, the improvement becomes saturated.
+Experiments on Text-Based LLMs and MLLMs with OCR results.
+To compare text-based LLMs and MLLMs performance on the MTVQA bench with OCR results, we adopt the Bytedance multilingual OCR API tool to extract multilingual text from images.
+For text-based LLMs, we choose GPT-4, and for MLLMs, we choose GPT-4V.
+As shown in Table  4, GPT-4V and OCR+GPT-4 have similar performance, but both are much lower than OCR+GPT-4V.
+We analyze the cases in detail and find differences in the challenges encountered by OCR+GPT-4 and GPT-4V.
+Since the performance of Latin languages is weaker than that of GPT4V(strong Latin text perception and comprehension abilities), OCR+GPT4 issues lie more in the lack of perception of visual elements and position.
+Many questions are dependent on visual elements and positional relationships in the image, resulting in that cannot be answered.
+GPT-4V issues lie more in the weakness of perception and comprehension of visual text, especially non-Latin text (AR/JA/KO/TH in Table 4), but GPT-4V can capture visual elements and positional relationships in images.
+More importantly, all three settings perform poorly and the comprehension of multilingual visual text remains a challenging case.
 5 Limitation
@@
 This prompts our future continual endeavor to ensure comprehensive representation across the linguistic spectrum.
-Additionally, the dataset currently provides a single canonical response per question, which may not fully capture the range of answers corresponding to different expressions of the same underlying semantics.
-Recognizing the multifaceted nature of the inquiries, future versions will aim to include a spectrum of plausible answers to reflect the varied perspectives inherent to each question.
 6 Conclusion
@@
 We believe that MTVQA is the first benchmark to provide fully manual annotations tailored to text-centric scenarios.
-The results obtained from closed-source, general-purpose and text-centric MLLMs on our MTVQA dataset indicate that there is still room for improving their performance in multilingual text-centric scenarios.
+The results obtained from closed-source, general-purpose, and text-centric MLLMs on our MTVQA dataset indicate that there is still room for improving their performance in multilingual text-centric scenarios.
 Although the MTVQA dataset has limitations, such as the underrepresentation of several lesser-spoken languages and single canonical answers, future updates will address these issues by expanding the multilingual scope and including a range of plausible answers.
 We are confident that this dataset can inspire researchers within the TEC-VQA community with new perspectives and ideas.
-Appendix A Appendix
-(a) German (DE) example
-(b) Italian (IT) example
-(c) French (FR) example
-(d) Japanese (JA) example
-(e) Korean (KO) example
-Figure 7: Multilingual VQA examples selected from five languages. The corresponding translations in English are in brackets.
-Table 3: Mean lengths of question-answer pairs in different languages of the training set and test set, using GPT-4o tokenizer.
-AR DE FR IT JA KO RU TH VI
-Training Set
-Question 8.29 8.72 9.73 12.05 12.43 11.74 11.56 11.35 11.21
-Answer 9.66 6.96 7.34 11.24 12.70 13.56 12.00 11.26 13.31
-Test Set
-Question 8.08 8.29 9.76 11.93 12.48 12.2 11.65 10.98 10.99
-Answer 7.95 6.67 6.61 11.04 12.55 13.61 14.42 12.08 13.00
-Refer to caption Refer to caption Refer to caption
-Refer to caption Refer to caption Refer to caption
-Refer to caption Refer to caption Refer to caption
-Figure 8: Word clouds showcasing top questions in various languages, tokenized via NLTK with the removal of stop words, punctuation, and digits.

### v4

@@
 In this work, we tackle multilingual TEC-VQA by introducing MTVQA, the first benchmark featuring high-quality human expert annotations across 9 diverse languages, consisting of 6,778 question-answer pairs across 2,116 images.
-Further, by comprehensively evaluating numerous state-of-the-art Multimodal Large Language Models (MLLMs), including Qwen2-VL, GPT-4o, GPT-4V, Claude3, and Gemini, on the MTVQA benchmark, it is evident that there is still a large room for performance improvement (Qwen2-VL scoring 30.9 versus 79.7 for human performance), underscoring the value of MTVQA.
-Additionally, we supply multilingual training data within the MTVQA dataset, demonstrating that straightforward fine-tuning with this data can substantially enhance multilingual TEC-VQA performance.
-We aspire that MTVQA will offer the research community fresh insights and stimulate further exploration in multilingual visual text comprehension.
+Further, by comprehensively evaluating numerous state-of-the-art Multimodal Large Language Models (MLLMs), including Qwen2.5-VL, InternVL-2.5, GPT-4o, GPT-4V, Claude3, and Gemini, on the MTVQA benchmark, it is evident that there is still a large room for performance improvement (InternVL-2.5 scoring 32.2 versus 79.7 for human performance), underscoring the value of MTVQA.
+By providing a dataset with nuanced multilingual annotations, MTVQA aims to set a new standard for benchmarks, fostering advancements in multilingual visual text comprehension.
+MTVQA is released at https://github.com/bytedance/MTVQA.
+MTVQA: Benchmarking Multilingual Text-Centric Visual Question Answering
+Jingqun Tang1, Qi Liu1, Yongjie Ye1, Jinghui Lu1, Shu Wei1, Chunhui Lin1, Wanqing Li1, Mohamad Fitri Faiz Bin Mahmood1, Hao Feng1, Zhen Zhao1, Yanjie Wang1, Yuliang Liu2, Hao Liu1 †, Xiang Bai2, Can Huang1 † 1ByteDance Inc.
+2Huazhong University of Science and Technology {tangjingqun, liuqi.nero, yeyongjie.ilz, lujinghui, haoliu.0128, can.huang}@bytedance.com, {ylliu, xbai}@hust.edu.cn
+footnotetext: Equal contribution. † Corresponding authors.
 1 Introduction
-In the era of burgeoning AI, especially in MLLMs [34, 1, 47, 44, 3, 39, 5, 27, 49, 12, 13, 18, 26, 43, 9, 11, 22, 24], Text-Centric Visual Question Answering (TEC-VQA[6, 42, 13, 12, 43, 26, 18] has served as a de facto gold proxy to evaluate AI models in the domain of text-centric scene understanding.
-Compared with general VQA [6, 30, 36, 42, 32, 31, 29, 52, 20, 4, 28, 40, 25, 15, 14, 23], TEC-VQA places greater emphasis on answering questions that require understanding visual textual information within images.
+In the era of burgeoning AI, especially in MLLMs (OpenAI, 2024; Team et al., 2023; Guo et al., 2025; Wang et al., 2025c; Anthropic, 2024; Bai et al., 2023; Lu et al., 2024a; Young et al., 2024; Chen et al., 2024b; Dong et al., 2024; Li et al., 2024; Liu et al., 2024a; Wang et al., 2025a), Text-Centric Visual Question Answering (TEC-VQA(Biten et al., 2019; Singh et al., 2019; Feng et al., 2023b, 2024; Tang et al., 2024; Zhao et al., 2024b, 2025b; Hu et al., 2024; Feng et al., 2025; Lu et al., 2024b, 2025) has served as a de facto gold proxy to evaluate AI models in the domain of text-centric scene understanding.
+Compared with general VQA, TEC-VQA places greater emphasis on answering questions that require understanding visual text information within images.
 It enables individuals without specialized expertise to access applications in text-centric visual environments.
-However, most advancements in TEC-VQA have predominantly concentrated on high-resource languages, e.g., English [6, 42, 30, 31], Chinese [37, 15], Japanese [41, 33] and etc., thus restricting the applicability of AI models to the global community, particularly populations speaking low-resource languages.
-To tackle the problem of language diversity, several seminal studies [38, 35, 7] in the general VQA field simply leverage translation engines to expand question-answer pairs from high-resource to low-resource languages.
+However, most advancements in TEC-VQA have predominantly concentrated on high-resource languages, e.g., English (Biten et al., 2019; Singh et al., 2019; Mathew et al., 2021, 2022; Wang et al., 2025b), Chinese (Qi et al., 2022; Gao et al., 2015; Fu et al., 2024), Japanese (Shimizu et al., 2018; Nguyen et al., 2023), and etc., thus restricting the applicability of AI models to the global community, particularly populations speaking low-resource languages.
+To tackle the problem of language diversity, several seminal studies (Raj Khan et al., 2021; Pfeiffer et al., 2022; Changpinyo et al., 2023) in the general VQA field simply leverage translation engines to expand question-answer pairs from high-resource to low-resource languages.
 However, this translation-based approach is not feasible for TEC-VQA, as it merely processes text in question-answer pairs, neglecting the critical visual text needed for answering based on images.
-Although the visual text can be recognized by an OCR (Optical Character Recognition) engine and then translated into the target language, this indirect process could lead to a “visual-textual misalignment” problem, due to nuanced meaning, contextual distortion, language bias, and question type diversity.
+Although the visual text can be recognized by an OCR (Optical Character Recognition) (Zhao et al., 2024a; Tang et al., 2022c, b, a, 2023; Liu et al., 2023) engine and then translated into the target language, this indirect process could lead to a “visual-textual misalignment” problem, due to nuanced meaning, contextual distortion, language bias, and question type diversity.
 Taking the second case in Fig.
@@
 In this work, to answer the question above, we establish MTVQA, a novel and high-quality multilingual TEC-VQA benchmark, where all images are collected from real-world and meticulously annotated by human experts in nine languages: Arabic (AR), Korean (KO), Japanese (JA), Thai (TH), Vietnamese (VI), Russian (RU), French (FR), German (DE), and Italian (IT).
-More concretely, to ensure the visual-textual alignment at most, the annotation process follows the raise-then-correct paradigm, where a group of human annotators raises several distinct questions, ranging from simple content extraction to text-related reasoning, and subsequently provides answers.
+More concretely, to ensure the visual-textual alignment at most, the annotation process follows the maker-checker paradigm, where a group of human annotators raises several distinct questions, ranging from simple content extraction to text-related reasoning, and subsequently provides answers.
 Another group then double-checks these QA pairs to ensure accuracy and consistency.
@@
 2, 6,678 training images and 21,829 question-answer pairs, as well as 2,116 test images and 6,778 question-answer pairs are obtained, covering more than 20 fine-grained scenarios from both documents and natural scenes, such as menus, logos, maps, bills, PPTs, research papers, and etc.
-To our knowledge, MTVQA is the first TEC-VQA dataset to provide native human annotations for multilingual text-rich scenarios, especially for low-source languages.
-Figure 3: Left: comparison of the overall performance of various MLLMs in the MTVQA benchmark. Right: comparison of the performance exhibited by MLLMs in the 9 languages of the MTVQA.
-We further investigate recent representative MLLMs in Fig.
-3, including GPT-4o, GPT-4V, Gemini, Qwen2-VL, etc., and their performance on MTVQA.
-As vividly revealed in Fig.
-3 (a), the general MLLM Qwen2-VL is the top performer, followed by closed-source GPT-4o and Claude3 Opus.
+To our knowledge, MTVQA is the first TEC-VQA dataset to provide native human annotations for multilingual text-rich scenarios, especially for low-resource languages.
+We further investigate recent representative MLLMs, including GPT-4o, GPT-4V, Gemini, Qwen2.5-VL, InternVL-2.5, etc., and their performance on MTVQA.
+As vividly revealed in Tab.
+2, the general MLLM InternVL-2.5 is the top performer, followed by Qwen2.5-VL and closed-source GPT-4o and Claude3 Opus.
 Other general MLLMs like QwenVL Max and QwenVL Plus show mid-tier performance, while text-centric MLLMs lag behind.
-Fig.
-3(b) shows closed-source MLLMs consistently outperforming others, especially in French (FR) and German (DE), while languages like Arabic (AR) and Thai (TH) witness lower scores across all categories.
 The results unequivocally demonstrate that opportunities for improvement persist within existing MLLMs when applied in multilingual text-rich scenarios.
@@
   • We establish a set of new multilingual TEC-VQA baselines for closed-source and general-purpose, text-centric MLLMs.
-  • MTVQA has been serving as a standard benchmark for evaluating multilingual visual text comprehension capabilities by some recently released state-of-the-art MLLMs such as Qwen2-VL [46] and InternVL 2 [11].
 2 Related Work
-Table 1: Comprehensive comparison on the benchmarks related to MTVQA. GPT4V denotes GPT-4V performance.
 Benchmark Scene Manual QA QA Language Visual Text Language GPT4V
@@
 MTVQA Multiple Text-rich 9 Languages 10 languages 22.0%
+Table 1: Comprehensive comparison on the benchmarks related to MTVQA. GPT4V denotes GPT-4V performance.
 2.1 MLLMs for Text-centric VQA
-Recent advancements in MLLMs [1, 47, 44, 3, 39, 5, 27, 49, 12, 13, 18, 26, 43, 9, 11, 22, 24, 51] have revolutionized VQA tasks, as demonstrated by the remarkable zero-shot performance of these models.
-Notably, the high generalizability of MLLMs, when explicitly trained on visual text understanding datasets and fine-tuned with instructions, has significantly enhanced their application in text-centric VQA scenarios [13, 12, 43, 26, 18].
-For example, LLaVAR [50], UniDoc [13], which extend LLaVA [25] into the realm of document understanding, pioneering the text-centric VQA of MLLMs by training them to predict texts and coordinates from document images.
-Furthermore, DocPedia [12] operates visual input in the frequency domain rather than in space, which enables higher input resolution without increasing the input sequence.
-Lately, mPLUG-DocOwl [48], Qwen-VL [5], and TextMonkey [26] leverage publicly available document-related VQA datasets to further enhance the text-centric VQA capabilities.
-Despite the promising results achieved by existing MLLMs in text-centric VQA tasks, their focus on high-resource languages such as English and Chinese has posed challenges in achieving reasonable performance for low-resource languages.
+Recent advancements in MLLMs have revolutionized VQA tasks, as demonstrated by the remarkable zero-shot performance of these models.
+Notably, the high generalizability of MLLMs, when explicitly trained on visual text understanding datasets and fine-tuned with instructions  (Lu et al., 2025; Fei et al., 2025; Ke et al., 2025; Qiu et al., 2025; Zhao et al., 2025a; Ouyang et al., 2024; Shen and Tang, 2024; Shen et al., 2025), has significantly enhanced their application in text-centric VQA scenarios (Feng et al., 2023b, a; Tang et al., 2024; Liu et al., 2024c; Hu et al., 2024).
+For example, LLaVAR (Zhang et al., 2023), UniDoc (Feng et al., 2023b), which extend LLaVA (Liu et al., 2024b) into the realm of document understanding, pioneering the text-centric VQA of MLLMs by training them to predict texts and coordinates from document images.
+Furthermore, DocPedia (Feng et al., 2023a) operates on visual input in the frequency domain rather than in space, which enables higher input resolution without increasing the input sequence.
+Lately, mPLUG-DocOwl (Ye et al., 2023), Qwen-VL (Bai et al., 2023), and TextMonkey (Liu et al., 2024c) leverage publicly available document-related VQA datasets to further enhance the text-centric VQA capabilities.
+What’s more, TextHarmony (Zhao et al., 2024b) unifies visual text generation and comprehension in a single model.
+Despite the promising results achieved by existing MLLMs in text-centric VQA tasks (He et al., 2025; Sun et al., 2025; Shan et al., 2024; Wang et al., 2025c), their focus on high-resource languages such as English and Chinese has posed challenges in achieving reasonable performance for low-resource languages.
 This is primarily due to the lack of data or benchmarks for these low-resource languages.
 2.2 Multilingual Text-centric VQA Benchmarks
-VQA has garnered significant attention in recent years, with numerous studies, datasets, and benchmarks being proposed to advance the field [6, 30, 36, 42, 32, 31, 29, 52, 20, 4, 28, 40, 25, 15, 14, 23].
-Many datasets have been created that encompass scene text of various domains, including natural images [6, 42], scanned documents [30, 31], book and movie covers [32].
-One notable limitation of these datasets is their predominant focus on English [6, 42, 30, 31] or other high-resource languages such as Chinese [37, 15] and Japanese [41, 33], which restricts the applicability of VQA systems for low-resource languages such as Thai and Vietnamese.
-There are some recent efforts toward extending VQA tasks to a broader range of languages [17, 35, 45, 7, 21, 38] by providing a multilingual VQA datasets.
+VQA has garnered significant attention in recent years, with numerous studies, datasets, and benchmarks being proposed to advance the field.
+Many datasets have been created that encompass scene text of various domains, including natural images (Biten et al., 2019; Singh et al., 2019), scanned documents (Mathew et al., 2021, 2022), books, and movie covers (Mishra et al., 2019).
+One notable limitation of these datasets is their predominant focus on English (Biten et al., 2019; Singh et al., 2019; Mathew et al., 2021, 2022) or other high-resource languages such as Chinese (Qi et al., 2022; Gao et al., 2015) and Japanese (Shimizu et al., 2018; Nguyen et al., 2023), which restricts the applicability of VQA systems for low-resource languages such as Thai and Vietnamese.
+There are some recent efforts toward extending VQA tasks to a broader range of languages (Gupta et al., 2020; Pfeiffer et al., 2022; Vivoli et al., 2022; Changpinyo et al., 2023; Li et al., 2023; Raj Khan et al., 2021) by providing multilingual VQA datasets.
 For example,  Gao et al.
-[15] created a free-form bilingual VQA dataset (FM-IQA) containing over 150,000 images and 310,000 freestyle Chinese question-answer pairs and English translations.
+(2015) created a free-form bilingual VQA dataset (FM-IQA) containing over 150,000 images and 310,000 freestyle Chinese question-answer pairs and English translations.
 Raj Khan et al.
-[38] developed a large-scale multilingual and code-mixed VQA dataset (MuCo-VQA) supporting five languages.
-Of more relevance are the works xGQA (7 languages) [35] and MaXM (7 languages) [7], which apply translation-based protocols to expand VQA data beyond English.
+(2021) developed a large-scale multilingual and code-mixed VQA dataset (MuCo-VQA) supporting five languages.
+Of more relevance are the works xGQA (7 languages) (Pfeiffer et al., 2022) and MaXM (7 languages) (Changpinyo et al., 2023), which apply translation-based protocols to expand VQA data beyond English.
 However, the translation-based multilingual VQA datasets inherently face issues, such as the “visual-textual misalignment” problem, where only the textual information in question-answer pairs is considered, while the visual text in images is overlooked.
-Additionally, the nuanced meaning and context are often distorted; language bias is introduced by machine translation models, and the coverage of certain question types is limited, as highlighted by [7].
+Additionally, the nuanced meaning and context are often distorted; language bias is introduced by machine translation models, and the coverage of certain question types is limited, as highlighted by Changpinyo et al.
+(2023).
 Moreover, none of the previous multilingual datasets focus on text-centric scenarios where multilingual text frequently occurs.
-Our benchmark, MTVQA, distinguishes itself by focusing on multilingual text-centric VQA scenarios using human expert annotations.
-It covers 9 languages, facilitating the training and evaluation of multilingual models in diverse linguistic contexts.
-Additionally, our dataset can gauge the VQA system’s ability for not only high-resource languages but also those that are typically underrepresented in current datasets [6, 42, 30, 31, 15].
 3 MTVQA Benchmark
@@
 The raw data collection aims to gather text-rich images from various scene text and document scenarios, ensuring diversity and quality.
-This includes images from publicly available datasets (e.g., ICDAR MLT19 [16]) and those sourced from the internet (i.e., Laion-OCR, which is filtered from Common Crawl [10]), such as menus, logos, maps, bills, PowerPoint slides (PPTs), research papers, etc, as shown in Fig.
+This includes images from publicly available datasets (e.g., ICDAR MLT19 (Gao et al., 2019)) and those sourced from the internet (i.e., Laion-OCR, which is filtered from Common Crawl (Crawl, 2024)), such as menus, logos, maps, bills, PowerPoint slides (PPTs), research papers, etc, as shown in Fig.
 2.
-The image collection process consists of two steps: extracting multilingual text using a multilingual OCR engine (Bytedance multilingual OCR API [^1] https://cv-api.bytedance.com/api/ocr-process/v1/multi_language_v2) and selecting by language types and the amount of text contained in the image.
+The image collection process consists of two steps: extracting multilingual text using a multilingual OCR engine (Bytedance multilingual OCR APIhttps://cv-api.bytedance.com/api/ocr-process/v1/multi_language_v2) and selecting by language types and the amount of text contained in the image.
 Approximately 30% of the overall image data is obtained from public datasets, with 20% sourced from the web and 50% from manual collection, respectively.
@@
 A total of 1,220 images from document scenarios and 876 images from natural scenarios are collected for the test set of the MTVQA benchmark.
-To ensure data content meets regulatory requirements, we subject them to a standardized data cleaning process.
+To ensure data content meets regulatory requirements, we subject it to a standardized data cleaning process.
 The image-cleaning processing pipeline involves a preliminary round of algorithm-driven review to identify and filter out unusable images.
@@
 One group is tasked with generating questions and providing answers based on the images, while the other group is responsible for evaluating and correcting the QA pairs.
-This division, known as the raise-then-correct paradigm, ensures a thorough and trustworthy evaluation of the text-rich image comprehension process.
+This division, known as the maker-checker paradigm, ensures a thorough and trustworthy evaluation of the text-rich image comprehension process.
 Moreover, the annotation results for each language underwent a 10% sampling inspection by a quality inspector to ensure adherence to standards.
@@
 The two-round annotation process is briefly illustrated in Fig.
-4, further details of which are elaborated in the subsequent sections.
+3, further details of which are elaborated in the subsequent sections.
 First Round Questioning and Answering.
@@
 Refer to caption Refer to caption Refer to caption
-Figure 5: Statistics of question and answer lengths of different languages aggregating training and test set, using GPT-4o tokenizer.
+Figure 4: Statistics of question and answer lengths of different languages aggregating training and test sets, using GPT-4o tokenizer.
 Refer to caption Refer to caption Refer to caption
@@
 Additionally, the benchmark showcases the vocabulary richness for each language through word clouds, as depicted in Fig.
-8, and the lengths of questions and answers are statistically analyzed using the GPT-4o tokenizer, as shown in Fig.
-5.
+7, and the lengths of questions and answers are statistically analyzed using the GPT-4o tokenizer, as shown in Fig.
+4.
 4 Experiments
@@
 To comprehensively assess MLLMs’ multilingual perception and comprehension capabilities, we select state-of-the-art MLLMs from three categories: open-source general MLLMs, open-source text-centric MLLMs, and closed-source MLLMs.
-Each category contains the following models: (1) General MLLMs: Qwen2-VL  [46], InternVL-V1.5 [8], InternLM-Xcomposer2-4KHD [11], Mini-Gemini-HD-34B [22], Llava-Next-34B [24], DeepSeek-VL [27], YI-VL-34B [49]; (2) Text-centric MLLMs: TextSquare [43], TextMonkey [26], mPLUG-DocOwl 1.5 [18], MiniCPM-V 2.0 [19]; (3) Closed-source MLLMs: GPT-4o [34], GPT-4V [1], Gemini Ultra [44], QwenVL Max [5], QwenVL Plus [5], Claude3 Opus [3], Claude3 Sonnet [3], and GLM-4V [2].
-For the closed-source MLLMs, we use the chat version through official APIs, while for the open-source MLLMs, we utilize the instruct versions available on the HuggingFace Model Hub.
-The open-source MLLMs’ model size varies from 7b to 76b.
-models GPT-4V GPT-4o Gemini Ultra Claude3 Opus Qwen2-VL 92B InternVL2 76B
-ANLS 29.6 36.9 31.7 35.4 38.7 29.1
-Table 2: Performance of the leading closed- and open-source MLLMs on the MTVQA benchmark.
-The best results of each language are bolded.
-The second best results are underlined.
-“Xcomposer-SFT” denotes instruction tuning to Xcomposer2-4KHD with MTVQA’s training set.
 AR DE FR IT JA KO RU TH VI Avg.
@@
 Closed-Source MLLMs
-GPT-4V [1] 11.5 31.5 40.4 32.3 11.5 16.7 10.3 15.0 28.9 22.0
-GPT-4o [34] 20.2 34.2 41.2 32.7 20.0 33.9 11.5 22.5 34.2 27.8
-GPT-4o mini [34] 16.9 33.0 41.2 32.1 18.5 27.4 11.5 19.9 29.1 25.5
-Gemini Ultra [44] 14.7 32.3 40.0 31.8 12.3 17.2 11.8 20.3 28.6 23.2
-QwenVL Max [5] 7.7 31.4 37.6 30.2 18.6 25.4 10.4 4.8 23.5 21.1
-QwenVL Plus [5] 4.8 28.8 33.7 27.1 12.8 19.9 9.4 5.6 18.1 17.8
-Claude3 Opus [3] 15.1 33.4 40.6 34.4 19.4 27.2 13.0 19.5 29.1 25.7
-Claude3 Sonnet [3] 10.5 28.9 35.6 31.8 13.9 22.2 11.0 15.2 20.8 21.1
-GLM-4V [2] 0.3 30.0 34.1 30.1 3.4 5.7 3.0 3.5 12.3 13.6
+GPT-4V (Achiam et al., 2023) 11.5 31.5 40.4 32.3 11.5 16.7 10.3 15.0 28.9 22.0
+GPT-4o (OpenAI, 2024) 20.2 34.2 41.2 32.7 20.0 33.9 11.5 22.5 34.2 27.8
+GPT-4o mini (OpenAI, 2024) 16.9 33.0 41.2 32.1 18.5 27.4 11.5 19.9 29.1 25.5
+Gemini Ultra (Team et al., 2023) 14.7 32.3 40.0 31.8 12.3 17.2 11.8 20.3 28.6 23.2
+QwenVL Max (Bai et al., 2023) 7.7 31.4 37.6 30.2 18.6 25.4 10.4 4.8 23.5 21.1
+QwenVL Plus (Bai et al., 2023) 4.8 28.8 33.7 27.1 12.8 19.9 9.4 5.6 18.1 17.8
+Claude3 Opus (Anthropic, 2024) 15.1 33.4 40.6 34.4 19.4 27.2 13.0 19.5 29.1 25.7
+Claude3 Sonnet (Anthropic, 2024) 10.5 28.9 35.6 31.8 13.9 22.2 11.0 15.2 20.8 21.1
+GLM-4V (AI, 2024) 0.3 30.0 34.1 30.1 3.4 5.7 3.0 3.5 12.3 13.6
 Open-Source MLLMs
-Qwen2-VL 72B [46] 20.7 36.5 44.1 42.8 21.6 37.4 15.6 17.7 41.6 30.9
-InternVL2 76B [1] 9.5 31.3 35.7 35.2 11.1 14.3 11.9 10.0 26.9 22.0
-InternVL-V1.5 [8] 3.4 27.1 31.4 27.1 9.9 9.0 4.9 8.7 12.4 14.9
-Mini-Gemini-HD-34B [22] 2.2 25.0 29.2 25.5 6.1 8.6 4.1 4.3 11.8 13.0
-Llava-Next-34B [24] 3.3 24.0 28.0 22.3 3.6 6.1 2.6 0.4 9.8 11.1
-DeepSeek-VL [27] 0.6 14.2 15.3 15.2 2.9 3.8 1.6 0.9 5.2 6.6
-YI-VL-34B [49] 1.7 13.5 15.7 12.1 4.8 5.2 0.8 3.5 4.1 6.8
-MiniCPM-V 2.0 [19] 1.3 12.7 14.9 17.0 3.7 5.6 2.2 2.2 6.8 7.4
-MiniCPM-V 2.5 [19] 6.1 29.6 35.7 26.0 12.1 13.1 5.3 12.6 15.3 17.3
-TextSquare [43] 3.7 27.0 30.8 26.7 3.2 7.2 6.7 5.2 12.4 13.6
-TextMonkey [26] 2.0 18.1 19.9 22.1 4.6 7.2 3.2 0.9 11.1 9.9
-mPLUG-DocOwl 1.5 [18] 1.0 13.9 14.9 18.2 2.9 5.0 2.0 0.9 6.4 7.2
-Xcomposer2-4KHD [11] 2.0 20.6 23.2 21.6 5.6 7.7 4.1 6.1 10.1 11.2
+Qwen2-VL 72B (Wang et al., 2024) 20.7 36.5 44.1 42.8 21.6 37.4 15.6 17.7 41.6 30.9
+Qwen2.5-VL 72B (Qwen, 2025) 17.2 34.2 38.5 35.2 18.4 29.6 13.9 9.5 41.6 28.7
+InternVL2 76B (Chen et al., 2023) 9.5 31.3 35.7 35.2 11.1 14.3 11.9 10.0 26.9 22.0
+InternVL2.5 78B  (Chen et al., 2024a) 15.9 39.0 45.6 42.9 21.1 33.9 12.2 23.8 41.5 32.2
+Mini-Gemini-HD-34B (Li et al., 2024) 2.2 25.0 29.2 25.5 6.1 8.6 4.1 4.3 11.8 13.0
+Llava-Next-34B (Liu et al., 2024a) 3.3 24.0 28.0 22.3 3.6 6.1 2.6 0.4 9.8 11.1
+DeepSeek-VL (Lu et al., 2024a) 0.6 14.2 15.3 15.2 2.9 3.8 1.6 0.9 5.2 6.6
+YI-VL-34B (Young et al., 2024) 1.7 13.5 15.7 12.1 4.8 5.2 0.8 3.5 4.1 6.8
+MiniCPM-V 2.0 (Hu et al., 2023) 1.3 12.7 14.9 17.0 3.7 5.6 2.2 2.2 6.8 7.4
+MiniCPM-V 2.5 (Hu et al., 2023) 6.1 29.6 35.7 26.0 12.1 13.1 5.3 12.6 15.3 17.3
+TextSquare (Tang et al., 2024) 3.7 27.0 30.8 26.7 3.2 7.2 6.7 5.2 12.4 13.6
+TextMonkey (Liu et al., 2024c) 2.0 18.1 19.9 22.1 4.6 7.2 3.2 0.9 11.1 9.9
+mPLUG-DocOwl 1.5 (Hu et al., 2024) 1.0 13.9 14.9 18.2 2.9 5.0 2.0 0.9 6.4 7.2
+Xcomposer2-4KHD (Dong et al., 2024) 2.0 20.6 23.2 21.6 5.6 7.7 4.1 6.1 10.1 11.2
 Xcomposer-SFT 11.8 31.7 37.4 29.3 14.5 12.9 5.8 13.9 20.2 19.7
+Table 2: Performance of the leading closed- and open-source MLLMs on the MTVQA benchmark.
+The best results of each language are bolded.
+The second-best results are underlined.
+“Xcomposer-SFT” denotes instruction tuning to Xcomposer2-4KHD with MTVQA’s training set.
 4.2 Implementation Details
@@
 This approach aims to make the answers as concise as possible.
-Besides, InternLM-Xcomposer2-4KHD  [11] is chosen as the base model for an instruction tuning experiment on the MTVQA training set.
+Besides, InternLM-Xcomposer2-4KHD  (Dong et al., 2024) is chosen as the base model for an instruction tuning experiment on the MTVQA training set.
 The instruction tuning process adheres to the default training settings specified by the source, with “HD-16” and completes one epoch of training on 8 NVIDIA-A100 GPUs within 2 hours.
+models GPT-4V GPT-4o Gemini Claude3 Qwen2-VL 72B InternVL2.5 78B
+ANLS 29.6 36.9 31.7 35.4 38.7 41.6
+Acc 22.0 27.8 23.2 25.7 30.9 32.2
+Table 3: Model performance comparison under ANLS and Acc metric.
+AR DE FR IT JA KO RU TH VI Avg.
+OCR+GPT-4 18.9 21.2 29.8 27.3 15.8 24 10.6 19 28.2 21.6
+OCR+GPT-4V 22.3 35.1 42.6 36.7 19.2 32.1 12.1 20.9 34.1 28.3
+GPT-4V 11.5 31.5 40.4 32.3 11.5 16.7 10.3 15 28.9 22.0
+Table 4: Comparison on applying OCR to GPT-4 and GPT-4V.
+4.3 Evaluation Results
+4.3.1 Evaluation metric
+To accurately assess whether the visual text that occurs in the answer is correct, we adopt Accuracy as the metric.
+The Accuracy metric measures the percentage of questions for which the predicted answer matches any of the target answers for the question.
+We also compare the models and human performance using the ANLS and Accuracy metrics.
+The models perform a little higher (5% - 10%) when using ANLS(Tab.3), but the rankings are unchanged.
+Since the ANLS metrics do not accurately reflect whether the correct textual content in the image is being presented, we chose the Accuracy metric.
+4.3.2 Human performance
+To establish a benchmark for human performance on MTVQA, 10 well-educated native speakers are randomly selected for each language.
+As shown in Table 2, the average human performance on MTVQA is around 80%, which is far beyond the performance of state-of-the-art MLLMs.
+It demonstrates the shortcomings and the huge room for improvement of the existing models in multilingual visual text perception and comprehension.
+4.3.3 Zero-shot evaluation
+[…]
+4.3.4 Instruction tuning
+[…]
+4.3.5 Experiments on Text-Based LLMs and MLLMs with OCR results
+To compare text-based LLMs and MLLMs’ performance on the MTVQA bench with OCR results, we adopt the Bytedance multilingual OCR API tool to extract multilingual text from images.
+For text-based LLMs, we choose GPT-4, and for MLLMs, we choose GPT-4V.
+As shown in Table  4, GPT-4V and OCR+GPT-4 have similar performance, but both are much lower than OCR+GPT-4V.
+We analyze the cases in detail and find differences in the challenges encountered by OCR+GPT-4 and GPT-4V.
+Since the performance of Latin languages is weaker than that of GPT4V(strong Latin text perception and comprehension abilities), OCR+GPT4 issues lie more in the lack of perception of visual elements and position.
+Many questions are dependent on visual elements and positional relationships in the image, resulting in which cannot be answered.
+GPT-4V issues lie more in the weakness of perception and comprehension of visual text, especially non-Latin text (AR/JA/KO/TH in Table 4), but GPT-4V can capture visual elements and positional relationships in images.
+More importantly, all three settings perform poorly, and the comprehension of multilingual visual text remains a challenging case.
+4.3.6 Failure Case Analysis
+We have compiled preliminary statistics on common errors: OCR failures (39%), reasoning gaps (34%), language bias (15%), and hallucinations (12%).
+For example, OCR Failure: Model misreads "café" as "cafe" in a French menu.
+Reasoning Gap: Fails to infer "9:00–16:00" as "business hours" in a Russian poster.
+5 Dataset Licenses
+The MTVQA dataset could be used publicly under the CC BY-NC 4.0 license.
+6 Acknowledgement
+This work is supported by the Natural Science Foundation of China (NSFC) NO. 62225603.
+7 Conclusion
+To attack the visual-textual misalignment issue in multilingual TEC-VQA, we introduce MTVQA, a new benchmark featuring high-quality human expert annotations in 9 diverse languages.
+We believe that MTVQA is the first benchmark to provide fully manual annotations tailored to text-centric scenarios.
+The results obtained from closed-source, general-purpose, and text-centric MLLMs on our MTVQA dataset indicate that there is still room for improving their performance in multilingual text-centric scenarios.
+Although the MTVQA dataset has limitations, such as the underrepresentation of several lesser-spoken languages, future updates will address these issues by expanding the multilingual scope and including a range of plausible answers.
+We are confident that this dataset can inspire researchers within the TEC-VQA community with new perspectives and ideas.
+8 Limitation
+The current version of the MTVQA dataset, while dialectically diverse, exhibits limitations in its language coverage.
+Despite encompassing a range of languages, it falls short of inclusivity, omitting numerous lesser-spoken languages.
+This prompts our future endeavor to ensure comprehensive representation across the linguistic spectrum.
+Appendix A More Experiments
 Table 5: Few-shot performance of GPT-4V on the MTVQA benchmark.
 In-context examples are randomly selected from the training set of MTVQA in the respective languages.
-“n-shot” represents the number of the selected in-context examples.
+“n-shot” represents the number of selected in-context examples.
 AR DE FR IT JA KO RU TH VI Avg.
@@
 eight-shot 15.5 34.8 42.1 35.6 13.4 17.5 12.1 18.2 32.8 24.7
-Table 4: Comparison on applying OCR to GPT-4 and GPT-4V
-AR DE FR IT JA KO RU TH VI Avg.
-OCR+GPT-4 18.9 21.2 29.8 27.3 15.8 24 10.6 19 28.2 21.6
-OCR+GPT-4V 22.3 35.1 42.6 36.7 19.2 32.1 12.1 20.9 34.1 28.3
-GPT-4V 11.5 31.5 40.4 32.3 11.5 16.7 10.3 15 28.9 22.0
-4.3 Evaluation Results
-Evaluation metric.
-To accurately assess whether the visual text that occurs in the answer is correct, we adopt Accuracy as the metric.
-The Accuracy metric measures the percentage of questions for which the predicted answer matches with any of the target answers for the question.
-We also compare the models and human performance using the ANLS and Accuracy metrics.
-The models perform a little higher (5% - 10%) when using ANLS, but the rankings are mostly unchanged.
-Since the ANLS metrics do not accurately reflect whether the correct textual content in the image is being presented, we chose the Accuracy metric.
-Human performance.
-To establish a benchmark for human performance on MTVQA, 10 well-educated native speakers are randomly selected for each language.
-As shown in Table 2, the average human performance on MTVQA is around 80%.
-[…]
-[…]
-Few-shot evaluation of GPT-4V.
+A.1 Few-shot evaluation of GPT-4V.
 Here, we compare the performance of GPT-4V on MTVQA under the few-shot settings.
 Specifically, we perform zero-shot, two-shot, five-shot, and eight-shot evaluations.
-We randomly select in-context examples from the train set in the respective languages and evaluate GPT-4V on the remaining instances.
+We randomly select in-context examples from the training set in the respective languages and evaluate GPT-4V on the remaining instances.
 As shown in Tab.
-3, compared to the zero-shot setting, GPT-4V’s performance has improved considerably under few-shot, highlighting its exceptional in-context learning ability in multilingual text comprehension contexts.
+5, compared to the zero-shot setting, GPT-4V’s performance has improved considerably under few-shot, highlighting its exceptional in-context learning ability in multilingual text comprehension contexts.
 Moreover, comparisons based on varying numbers of in-context samples reveal that augmenting the in-context samples can aid in further enhancement; however, after reaching a certain volume, the improvement becomes saturated.
-Experiments on Text-Based LLMs and MLLMs with OCR results.
-To compare text-based LLMs and MLLMs performance on the MTVQA bench with OCR results, we adopt the Bytedance multilingual OCR API tool to extract multilingual text from images.
-For text-based LLMs, we choose GPT-4, and for MLLMs, we choose GPT-4V.
-As shown in Table  4, GPT-4V and OCR+GPT-4 have similar performance, but both are much lower than OCR+GPT-4V.
-We analyze the cases in detail and find differences in the challenges encountered by OCR+GPT-4 and GPT-4V.
-Since the performance of Latin languages is weaker than that of GPT4V(strong Latin text perception and comprehension abilities), OCR+GPT4 issues lie more in the lack of perception of visual elements and position.
-Many questions are dependent on visual elements and positional relationships in the image, resulting in that cannot be answered.
-GPT-4V issues lie more in the weakness of perception and comprehension of visual text, especially non-Latin text (AR/JA/KO/TH in Table 4), but GPT-4V can capture visual elements and positional relationships in images.
-More importantly, all three settings perform poorly and the comprehension of multilingual visual text remains a challenging case.
-5 Limitation
-The current version of the MTVQA dataset, while dialectically diverse, exhibits limitations in its language coverage.
-Despite encompassing a range of languages, it falls short of inclusivity, omitting numerous lesser-spoken languages.
-This prompts our future continual endeavor to ensure comprehensive representation across the linguistic spectrum.
-6 Conclusion
-To attack the visual-textual misalignment issue in multilingual TEC-VQA, we introduce MTVQA, a new benchmark featuring high-quality human expert annotations in 9 diverse languages.
-We believe that MTVQA is the first benchmark to provide fully manual annotations tailored to text-centric scenarios.
-The results obtained from closed-source, general-purpose, and text-centric MLLMs on our MTVQA dataset indicate that there is still room for improving their performance in multilingual text-centric scenarios.
-Although the MTVQA dataset has limitations, such as the underrepresentation of several lesser-spoken languages and single canonical answers, future updates will address these issues by expanding the multilingual scope and including a range of plausible answers.
-We are confident that this dataset can inspire researchers within the TEC-VQA community with new perspectives and ideas.
+A.2 MLLM Performance Changes When the Question is Asked in English.
+We translate the questions into English using GPT-4 and then perform a test on GPT4V.
+The result is shown in Table  6.
+The GPT4V performance is robust when the question is asked in English and the original language.
+Table 6: GPT4V performance changes when the question is asked in English
+AR DE FR IT JA KO RU TH VI Average
+English Questions 11.9 31.8 39.0 32.1 12.0 15.8 10.1 18.1 28.5 22.1
+Original Questions 11.5 31.5 40.4 32.3 11.5 16.7 10.3 15.0 28.9 22.0
+A.3 The Exploration of Additional Metrics
+We compare ANLS (soft metric) and Accuracy, and we choose Accuracy to reflect the consistency of the answer and the visual text content because accurate answers (e.g., price, date) are critical.
+We also pilot LAVE but observe low consistency (76.8%) with human judgments for non-Latin languages (Table 2), due to the lower low-resource linguistic performance of LLMs.
+Appendix B More Visualizations
+(a) German (DE) example
+(b) Italian (IT) example
+(c) French (FR) example
+(d) Japanese (JA) example
+(e) Korean (KO) example
+Figure 6: Multilingual VQA examples selected from five languages. The corresponding English translations are in brackets.
+Table 7: Mean lengths of question-answer pairs in different languages of the training set and test set, using GPT-4o tokenizer.
+AR DE FR IT JA KO RU TH VI
+Training Set
+Question 8.29 8.72 9.73 12.05 12.43 11.74 11.56 11.35 11.21
+Answer 9.66 6.96 7.34 11.24 12.70 13.56 12.00 11.26 13.31
+Test Set
+Question 8.08 8.29 9.76 11.93 12.48 12.2 11.65 10.98 10.99
+Answer 7.95 6.67 6.61 11.04 12.55 13.61 14.42 12.08 13.00
+Refer to caption Refer to caption Refer to caption
+Refer to caption Refer to caption Refer to caption
+Refer to caption Refer to caption Refer to caption
+Figure 7: Word clouds showcasing top answers in various languages, tokenized via NLTK with removing stop words, punctuation, and digits.